### v2

# TabArena: A Living Benchmark for Machine Learning on Tabular Data

Abstract
With the growing popularity of deep learning and foundation models for tabular data, the need for standardized and reliable benchmarks is higher than ever. However, current benchmarks are static. Their design is not updated even if flaws are discovered, model versions are updated, or new models are released. To address this, we introduce TabArena, the first continuously maintained living tabular benchmarking system. To launch TabArena, we manually curate a representative collection of datasets and well-implemented models, conduct a large-scale benchmarking study to initialize a public leaderboard, and assemble a team of experienced maintainers. Our results highlight the influence of validation method and ensembling of hyperparameter configurations to benchmark models at their full potential. While gradient-boosted trees are still strong contenders on practical tabular datasets, we observe that deep learning methods have caught up under larger time budgets with ensembling. At the same time, foundation models excel on smaller datasets. Finally, we show that ensembles across models advance the state-of-the-art in tabular machine learning and investigate the contributions of individual models. We launch TabArena with a public leaderboard, reproducible code, and maintenance protocols to create a living benchmark available at https://tabarena.ai.
Figure 1: TabArena-v0.1 Leaderboard. We evaluate models under default parameters, tuning, and weighted ensembling [1] of hyperparameters. Since TabICL and TabPFNv2 are not applicable to all datasets, we evaluate them on subsets of the benchmark in Figure 4 .
1 Introduction
Benchmarking tabular machine learning models is an arduous and error-prone process. With the rise of deep learning and foundation models for tabular data [2, 3, 4, 5, 6], benchmarking has become even more challenging for researchers and practitioners alike. While several benchmarks have been proposed in recent years, there is increasing skepticism towards data curation and evaluation protocols utilized [7, 8, 9]. Most importantly, many datasets used in benchmarks are outdated, contain problematic licenses, do not represent real tabular data tasks, or are biased through data leaks [7, 10].
Despite the growing awareness of such issues, benchmarks are rarely maintained after publication. Issues uncovered in follow-up studies are not addressed, and baselines for the state-of-the-art are stuck in time. Consequently, follow-up benchmarks reproduce shortcomings of prior work and do not compare to the actual state-of-the-art [11].
To address these issues, we argue for a paradigm shift from the currently used static benchmarks to a sustainable, living benchmarking system treated as software that is versioned, professionally maintained, and gradually improved by the community as an open-source project. With this goal in mind, we introduce TabArena.
TabArena is a living benchmarking system that makes benchmarking tabular machine learning models a reliable experience. TabArena is realized through our four main contributions:
  • We investigate $1053$ datasets used in tabular data research and carefully, manually curate a set of $51$ datasets out of these, representing real-world tabular data tasks.
[…]
  • We instantiate a public leaderboard available on tabarena.ai, release precomputed results for fast comparisons, and provide reproducible code to benchmark new methods.
  • We assemble a team of maintainers from different institutions with experience in maintaining open-source initiatives to keep the living benchmark up to date.
In this paper, we detail our curation protocols for building a sophisticated living benchmark and investigate the results of TabArena version 0.1, representing our initialization of the leaderboard.
TabArena-v0.1 Focus Statement. TabArena focuses on evaluating predictive machine learning models for tabular data. Our long-term vision is to make TabArena representative for all use cases of tabular data. For TabArena-v0.1, we initialize the benchmark focusing on the most common type of tabular machine learning problem: Tabular classification and regression for independent and identically distributed (IID) data, spanning the small to medium data regime. We explicitly leave for future work use cases such as non-IID data (e.g., temporal dependencies, subject groups, or distribution shifts); few-shot predictions, or very small data (e.g., less than 500 training samples); large data (e.g., more than 250 000250000250\,000250 000 training samples); or other machine learning tasks such as clustering, anomaly detection, subgroup discovery, or survival analysis.
Our results demonstrate: (A) The best performance for individual tabular machine learning models is generally achieved by post-hoc ensembling tuned hyperparameter configurations; (B) With tuning and ensembling, the best deep learning methods are equal to or better than gradient-boosted decision trees; (C) Tabular foundation models dominate for small data with strong in-context learning performance even without tuning; (D) Ensemble pipelines leveraging various models are the state-of-the-art for tabular data, but the best individual models do not contribute equally to the ensemble pipeline.
2 TabArena
TabArena is a living benchmark because of our protocols, which govern the curation of: (2.1) models and hyperparameter optimization; (2.2) datasets; (2.3) and evaluation design. Through continuous application and refinement of these protocols, we ensure TabArena remains current and maintained.
2.1 Models and Hyperparameter Optimization Protocol
TabArena is implemented as an extendable platform to support a wide range of machine learning models on tabular data. For instantiating TabArena, we curate $14$ state-of-the-art (foundation) models and two simple baselines. TabArena is created as a platform to benchmark each model to its full potential. Therefore, every included model represents a well-known baseline or was implemented in dialogue with the authors. Furthermore, we only run models on datasets within their restrictions to represent them fairly. This only affects TabPFNv2, which is restricted to datasets with up to $10,000$ training samples, $500$ features, and $10$ classes for classification tasks, and TabICL, which is constrained to classification tasks with up to $100,000$ training samples and $500$ features.
TabArena models are powered by three components: (1) implementation in a well-tested modeling framework used in real-world applications; (2) curated hyperparameter optimization protocols; (3) improved validation and ensembling strategies, including ensembling over instances of a single model class. Table 1 provides an overview of the models benchmarked in TabArena-v0.1.
Table 1: TabArena-v0.1 Models. We show all models included in our initialization of TabArena, the source of the search space, and a short version of the name. Moreover, we specify the model types: tree-based ([Uncaptioned image]), neural network ([Uncaptioned image]), pretrained foundation models ([Uncaptioned image]), and baseline ([Uncaptioned image]).
Model Short Name Search Space Type
Random Forests [12] RandomForest Prior Work + Us [Uncaptioned image]
Extremely Randomized Trees [13] ExtraTrees Prior Work + Us [Uncaptioned image]
XGBoost [14] XGBoost Prior Work + Us [Uncaptioned image]
LightGBM [15] LightGBM Prior Work + Us [Uncaptioned image]
CatBoost [16] CatBoost Prior Work + Us [Uncaptioned image]
Explainable Boosting Machine [17, 18] EBM Authors [Uncaptioned image]
FastAI MLP  [19] FastaiMLP Authors [Uncaptioned image]
Torch MLP  [19] TorchMLP Authors [Uncaptioned image]
RealMLP [20] RealMLP Authors [Uncaptioned image]
[…]
ModernNCA [21] ModernNCA Authors [Uncaptioned image]
TabPFNv2 [5] TabPFNv2 Authors [Uncaptioned image]
TabICL [22] TabICL - [Uncaptioned image]
TabDPT [23] TabDPT - [Uncaptioned image]
Linear / Logistic Regression Linear TabRepo [Uncaptioned image]
K-Nearest Neighbors KNN TabRepo [Uncaptioned image]
Implementation Framework.  For implementing models, we rely on functionalities from AutoGluon [19], an established machine learning framework used in practical applications. Each model is implemented within the standardized AbstractModel framework, which aligns with the scikit-learn [24] API, and includes: (1) model-agnostic preprocessing, (2) support for (inner) cross-validation with ensembling, (3) hyperparameter optimization, (4) evaluation metrics, (5) fold-wise training parallelization, (6) (customizable) model-specific preprocessing pipeline, (7) (customizable) early stopping and validation logic, and (8) unit tests. As a result, any model implemented in TabArena can be readily deployed for real-world use cases or within predictive machine learning systems. Moreover, the pipeline logic encompassing models within TabArena is implemented in a tested framework regularly used in real-world applications. Section C.1 summarizes further implementation details, and Section D.2 includes a detailed protocol for contributing models.
Cross-validation and Ensembles. As can be seen in various Kaggle competitions and academic studies, cf. [25, 8, 26, 20], for most datasets, peak performance requires ensembling strategies. Therefore, we default to using $8$-fold cross-validation ensembles [27] within TabArena. For all foundation models, we refit on training and validation data instead of using cross-validation ensembles, following recommendations from the authors of TabPFN and TabICL. In addition, we evaluate each tunable model using post-hoc ensembling [1] of different hyperparameter configurations.
Hyperparameter Optimization.  For each model, we curate a strong hyperparameter search space; for full details, see Appendix C.2. Where possible, we started with the search spaces from the original paper and finalized them in dialogue with the models’ authors. Otherwise, we curated search spaces from prior work. We evaluate $1$ default and a fixed set of $200$ randomly-sampled hyperparameter configurations for all models, except for TabICL and TabDPT. TabICL and TabDPT do not specify hyperparameter optimization (HPO) per the original paper and implementation; thus, we restrict ourselves to evaluating only their in-context learning performance. Each hyperparameter configuration is validated using $8$-fold (inner) cross-validation. We use the score of this (inner) cross-validation to select the best hyperparameter configuration.
For practical reasons, we restrict the time to evaluate one configuration on one train split of a dataset to $1$ hour. We analyze the impact of this time limit in Section A.2.
In a living benchmark, we expect users with different hardware to submit to the leaderboard. Thus, we do not constrain the hardware used to evaluate a configuration. We log the hardware used during benchmarking to enable analysis of the impact of computing power. Our recommended hardware for evaluating a configuration is 8 CPU cores, 32 GB of RAM, and 100 GB of disk space. Furthermore, we recommend using 1 GPU with 48 GB VRAM for GPU-scaling models, such as foundation models.
2.2 Datasets Protocol
Table 2: Comparison of Tabular Benchmarks. We systematically compare prior tabular benchmark studies across six characteristics. Inner and outer splits: the number of splits used for inner or outer validation: $1$ for holdout validation; - if the benchmark does not specify; and any other number specifies the total number of splits from (repeated) cross-validation. If a set is given, the benchmark uses different splits for different datasets. Ensembling: Whether the benchmark studies ensembling of configurations for individual models (), uses any other ensembling ((✓)), or uses no ensembling at all (). Manual Curation: Whether the benchmark filters datasets based on criteria beyond simple automation, () or not (). Datasets remaining: the number of datasets remaining after filtering by our criteria. Results available: Whether the benchmark shares no re-usable results (), only metric results ((✓)), or metric results and predictions (). HPO Limit: How hyperparameter optimization was limited in the number of configurations and/or hours.
#splits Manual #datasets Results HPO Limit
Benchmark inner outer Ensembling curation remaining available #confs. #hours
[28, 29] 1 10 9/72 (✓) 1 -
[30] 1 1 (✓) 1/11 100 6
[31] 1 $\{1,3\}$ (✓) 1/11 1000 -
[32] 1 $\{1,2,3,5\}$ 12/47 (✓) 400 -
[33] 1 10 13/196 (✓) 30 10
[34] {1, 3, 10} {1, 10, 100} 8/35 (✓) {-, 500} -
[35] - 10 (✓) 15/104 (✓) - 4
[…]
[8] 10 1 (✓) 1/10 100 -
[20] 1 10 (✓) 10/118 50 -
[36] 1 1 39/300 (✓) 100 -
[10] 1 1 (✓) 0/8 (✓) 100 -
[37] 8 3 19/200 200 200
TabArena (Ours) 8 {9, 30} 51/51 200 200
Many existing benchmarks were curated using semi-automated procedures to collect datasets according to simple characteristics, often to obtain as many datasets as possible. In contrast, we reject the notion of automatically collecting datasets without any sanity check. Instead, we focus on carefully, manually curating a representative collection of datasets. Although some previous benchmarks manually curated data, most of the included datasets do not meet the criteria of IID tabular datasets, as seen in Table 2. Our work indicates a turning point with the most extensive and conscientious manual curation effort for machine learning on IID tabular data so far.
We define criteria for data selection according to our focus statement and filter 1053 datasets used in 13 prior benchmarks accordingly. Figure 2 describes our selection process. Notably, only the deduplication step and size filters can be automated. Faithfully applying the other criteria requires manual human effort per-dataset, demonstrating the downsides of automated data curation procedures.
Figure 2: Data Curation Results. The figure shows why and how many datasets we filter based on our criteria. We filter datasets that are duplicates, not from a tabular domain, not a real predictive task, tiny, have quality or license issues, and are not IID.
Dataset Selection Criteria.  We comprehensively describe our selection criteria in Section B.1. In short, we selected datasets that fulfilled the following requirements: (1) The dataset and its predictive machine learning task are unique within our benchmark; (2) The dataset is IID, that is, a random split is appropriate for the underlying original task; (3) The dataset is not from a non-tabular modality, such as images, where it is unclear whether tabular machine learning is a reasonable alternative to domain-specific methods; (4) The dataset stems from a real random distribution, and is not generated, e.g., from a deterministic function; (5) The dataset was published explicitly for a predictive modeling task in a real-world application; (6) The dataset is small-to-medium-sized, i.e., it has at least $500$ and at most 250 000250000250\,000250 000 train samples; (7) We can use a version of the dataset without pre-applied problematic preprocessing, such as irreversible data leaks; (8) The dataset was originally published with a license allowing for scientific usage; (9) The dataset and its structured metadata can be automatically downloaded via a public API, or we are allowed to upload the dataset to a public API; (10) The dataset and its predictive task do not raise (obvious) ethical concerns.
As several criteria involve subjective judgment and human interpretation, we publicly share our curation insights at tabarena.ai/dataset-curation, including per-dataset notes detailing our observations, identified characteristics, and final assessments. To enhance the quality of TabArena’s datasets, we actively encourage the community to review and critique our evaluations.
Figure 3: Characteristics of Datasets in TabArena. On the left, we show the number of datasets per task type, license, source of the dataset, and age group. On the right, we show the number of features (columns) and samples (rows), as well as the percentage of categorical features per dataset.
Figure 3 summarizes the characteristics of our datasets; we share per-dataset details in Section B.2. Throughout our curation process, we noticed, among others, two trends. First, across benchmarks, the versions of datasets were often inconsistent. Benchmarks used the same name and source for a dataset, but with different preprocessing, features, or even targets. Thus, dataset-specific performance comparisons across benchmarks were often invalid. Second, the number of datasets that are truly suitable for benchmarking IID tabular data approaches is surprisingly small after carefully inspecting the tasks represented by datasets. We share more noteworthy observations in Section B.3.
Call for Data Contributions.TabArena is a living benchmark, and the datasets we curated are a significant part of this living, continuously maintained system. As mentioned above, we invite the community to scrutinize our curation efforts. Moreover, to keep TabArena up-to-date, we implore the broader scientific community to share their tabular data publicly. We are also actively looking for contributions of data outside of TabArena-v0.1’s focus, i.e., data that is non-IID, tiny, or large, as well as tasks where extracing features from other modalities is still considered a reasonable approach. Likewise, we invite the community to curate additional existing datasets. We provide a detailed protocol to contribute a new dataset to TabArena in Section D.3.
2.3 Evaluation Design Protocol
The TabArena leaderboard aims to assess the state-of-the-art for predictive machine learning. With this in mind, we design the evaluation to produce a reliable and representative leaderboard. To guard against randomness in the data and method significantly impacting our conclusions, we repeat our experiments per dataset. We employ a dataset-specific repetition strategy using more repeats for smaller datasets and fewer for larger datasets, because the significance of randomness decreases with dataset size. The strategy is as follows: (I) for datasets with less than 2 50025002\,5002 500 samples, we use $10$ times repeated $3$-fold outer cross-validation; (II) for all other datasets, we use $3$ repeats.
We evaluate models using the Elo rating system [38], following Chatbot Arena [39]. Elo is a pairwise comparison-based rating system where each model’s rating predicts its expected win probability against others, with a 400-point Elo gap corresponding to a 10 to 1 (91%) expected win rate. We calibrate 1000 Elo to the performance of our default random forest configuration across all figures, and perform 100 rounds of bootstrapping to obtain 95% confidence intervals, as in [39]. In our main results, Elo scores are computed using ROC AUC for binary classification, log-loss for multiclass classification, and RMSE for regression. The leaderboard offers just one view of the results; we also track additional aspects, such as alternative metrics, training-, and inference-efficiency, to enable diverse analyses. Furthermore, we provide scripts as part of TabArena to generate the leaderboard, evaluation plots, and inspect the results.
TabArena Reference Pipeline. Following the recommendations of [11], we include a reference pipeline in our benchmark. This reference pipeline is applied independently of the tuning protocol and constraints we constructed for models within TabArena. The reference pipeline aims to represent the performance easily achievable by a practitioner on a dataset.
We select the predictive machine learning system AutoGluon [19] (version 1.3, with the best_quality preset and $4$ hours for training) as the first official TabArena reference pipeline.
Additional Metadata.  Next to the main results, we save an extensive amount of additional metadata to enable future research and deeper model studies. We save the training time, inference time, precomputed results for various metrics, hyperparameters, hardware specification, and model predictions. We save the validation and test predictions of the final model, and of all models trained per-fold during inner cross-validation.
Living Benchmark.TabArena marks the start of an open-source initiative towards a continuously updated, collectively shared assessment of the state-of-the-art in tabular machine learning. Therefore, as one of the most crucial parts of TabArena, we define a protocol for researchers and practitioners to submit models to the live leaderboard. The protocol is detailed in Section D.4
3 Results
To initialize the leaderboard, we run $16$ curated models (Section 2.1) on $51$ curated datasets (Section 2.2) within TabArena’s evaluation design (Section 2.3). Likewise, we evaluate AutoGluon, the reference pipeline, on all datasets. We run TabM, ModernNCA, and the foundation models on GPU and all other models, as well as AutoGluon, on CPU. We give all runs $32$ GB of RAM and $100$ GB of disk space. We perform CPU runs in the cloud via Amazon Web Services on an M6i.2xlarge EC2 instance (eight cores Intel Xeon CPU). We perform GPU runs on an NVIDIA L40S with 48 GB VRAM and eight cores of an AMD EPYC 9334 CPU. Based on our logs, the compute time for all our experiments without parallelization, including overhead such as scheduling, is ${\sim}15$ wall-clock years.
3.1 Assessing Peak Performance: The TabArena-v0.1 Leaderboard
Figure 1 shows our main results, the leaderboard of TabArena-v0.1, which includes the performance of methods with default parameters, hyperparameter tuning, and weighted post-hoc ensembling [1] of different hyperparameter configurations. Furthermore, we present one leaderboard encompassing all models in Section A.1, where we impute missing datasets, representing the live leaderboard available on tabarena.ai. To provide a wide range of perspectives, we also share additional versions of the leaderboard in Section A.1, such as using alternative metrics or task-wise dataset subsets. Finally, we analyze the statistical significance of our comparison in Section A.5.
Post-Hoc Ensembled Deep Learning Models Dominate the Leaderboard. In line with previous work [33], CatBoost is ranked first in the conventional tuning regime (Figure 1). However, the TabArena-v0.1 leaderboard reveals that after post-hoc ensembling, neural networks are the strongest single models on average in TabArena-v0.1. Our results show that the peak performance of models is misrepresented unless post-hoc ensembling is used. This is evident in the observation that the top three models in our leaderboard (TabM, LightGBM, RealMLP; see Figure 1) would all be worse than the actual fourth-best model (CatBoost) without post-hoc ensembling. While practitioners must trade off the increased predictive performance with the increased inference cost (see Figure 5) [40], peak performance requires post-hoc ensembling. Furthermore, this shows that compared to existing benchmarks, TabArena is closer to correctly representing the currently achievable peak performance.
Tabular Foundation Models Lead on Small Datasets. As most foundation models are currently limited in their applicability, we evaluate them separately in Figure 4, presenting the leaderboard after restricting the datasets to the constraints of TabPFNv2 or TabICL. TabPFNv2 outperforms related approaches by a large margin, establishing tabular foundation models as the go-to solution for datasets within their constraints. Moreover, TabPFNv2 with tuning and post-hoc ensembling again outperforms AutoGluon, confirming the results by Hollmann et al. [5]. We expect tabular foundation models with wider applicability to be released and included in future versions of TabArena.
Figure 4: Leaderboard for TabPFNv2-compatible (left) and TabICL-compatible (right) datasets. For TabPFNv2, we obtain $33$ datasets ($\leq$ 10K training samples, $\leq 500$ features). For TabICL, we obtain $36$ classification datasets ($\leq$ 100K, $\leq 500$). Everything but the datasets is identical to Figure 1.
3.2 Holistic Benchmarking of Peak Performance with TabArena
In this subsection, we demonstrate how the design choices behind TabArena enable users to assess peak performance and evaluate the utility of models in ensembling pipelines.
Abandoning the Holdout Validation Fallacy.  The rich metadata we save for TabArena allows us to study what-if cases for our evaluation design. Figure 5 shows that when using holdout validation instead of cross-validation, all models are greatly underestimated, and performance rankings are biased in favor of models that already use ensembling. This aligns with prior work [11] and demonstrates the importance of nested cross-validation in our benchmark design.
Figure 5: (Left) Model Efficiency. Median times across TabPFNv2- and TabICL-compatible datasets. Training times include the time for cross-validation. (Right) Holdout vs. Cross-Validation. Predictive performance of a model with tuning and ensembling when using holdout or cross-validation.
Reducing Time and Hardware Constraints for Accurate Benchmarking.  While TabArena enables a more accurate assessment of peak performance, it does so at a non-negligible computational cost. We hypothesize that prior benchmarks did not assess peak performance mainly due to time and hardware constraints. Thus, we share our result artifact to reduce time and hardware constraints for accurate benchmarking. As a result, users can cheaply compare novel models, and simulated ensembles thereof, to existing models and ensemble pipelines (see Section D.1).
Furthermore, we introduce and continually maintain TabArena-Lite, a subset of TabArena that consists of all datasets with one outer fold (see Section A.4). We intend TabArena-Lite to be used in academic studies and find any novel model that outperforms other models on at least one dataset, even if it is not the state-of-the-art on average, worthy of publication.
Ensembles Across Models and Individual Contributions.  While our public leaderboard summarizes model capabilities in one metric, we want to raise awareness that this is a limited perspective, wrongly suggesting that lower-ranked models are unnecessary. In Figure 6 (left), we show that a simulated ensembling pipeline using all models in TabArena outperforms all individual models and AutoGluon, further advancing the state-of-the-art. At the same time, certain individual models are more important than others for the state-of-the-art. In Figure 6 (right), we show the average weights of different models in the TabArena ensemble. Notably, models with the highest performance on the leaderboard are not necessarily the ones with the highest weights, likely because the ensemble construction favors models that perform better on the validation than test data [41, 42].
A Sober Look at the “GBDT vs. Deep Learning” Debate. Prior benchmarks dedicated an enormous amount of work toward debating whether gradient-based decision trees (GBDTs) or deep learning is better across benchmarks and model studies [31, 32, 33]. We argue that the battle between GBDTs and deep learning is a false dichotomy, as both model families contribute to ensembles that strongly outperform individual model families (Figure 6). Thus, we posit that it is more important to find new models that work well in ensembles than models that beat GBDTs. TabArena enables such research by allowing users to simulate ensemble performance per and across models based on our precomputed result artifacts. Finally, to highlight, we perceive the most common reasons for practitioners to rely on GBDTs instead of deep learning to be the insufficient code quality and maintenance of deep learning methods compared to GBDT frameworks. TabArena also tackles this problem by wrapping existing deep learning methods in an easy-to-use interface, see Section D.5.
Figure 6: (Left) TabArena Ensemble. Simulated performance of an ensemble using all models in TabArena compared to the leaderboard from Figure 1. (Right) Individual Contributions. Contributions of models to the TabArena ensemble in terms of average weight across all datasets.
4 Related Work
Many tabular benchmarks have been proposed over the recent years, including OpenML-CC18 [28], PMLB(Mini) [43, 44, 45], the Grinsztajn et al. [32] benchmark, TabZilla [33], the AutoML Benchmark [35], TabRepo [37], Tschalzev et al. [8], TabRed [10], the [46] benchmark, and TALENT [47]. In addition, many studies used their own, alternative evaluation frameworks [30, 20, 48, 26, 9, 5]. This shows that the community does not have a consistent understanding of how to benchmark tabular models and, more importantly, no platform to align and collectively improve benchmarking efforts. TabArena aims to provide both.
Recent studies revealed further flaws in tabular benchmarking: datasets are often outdated [7], non-IID data is used inappropriately [10], many datasets are not originally tabular [10], and inappropriate evaluation protocols are frequent [8, 11]. TabArena addresses these concerns through extensive data curation efforts and evaluation in sophisticated ensembling pipelines. As seen in Table 2, TabArena consolidates the individually outstanding aspects of previous benchmarks.
While some existing tabular benchmarks provide guidelines to add datasets or models, none include maintenance protocols or have been updated to address issues discovered in data and evaluation quality. TabArena is the first living tabular benchmark with active maintenance protocols and a public leaderboard for tabular data. Our work is inspired by benchmarking efforts from other domains, such as ChatBot Arena [39] or LiveBench [49] for LLMs and GIFT-Eval [50] for time-series forecasting.
5 Conclusion
[…]
Limitations and Societal Impact.  As we envision a living benchmarking system that will evolve over time, some limitations can be seen as future work, while others stem from fundamental trade-offs in benchmark design choices. Our (current) limitations are: (1) We use a fixed set of 200 random hyperparameter configurations to enable the study of ensemble pipelines. This prevents analyzing the variance of random hyperparameter choices [51] and studying more advanced hyperparameter optimization strategies. (2) We use a time limit per configuration; thus, our results depend on the hardware used in edge cases where the time limit is reached. Using different hardware across models (and in the future across users) reduces the comparability in such cases. (3) Our strict selection criteria for datasets makes TabArena more representative for real-world use-cases, but reduce the number of datasets drastically. We emphasize the need to work with the community to curate a more representative, high-quality collection of datasets with more diversity and statistical power. (4) Lastly, we assess predictive performance without feature engineering on top of the existing dataset state. Feature engineering could further boost predictive performance and change the model ranking.
Our work has a broader positive societal impact by improving the trustworthiness and reliability of academic benchmarks. Moreover, TabArena provides practical guidance to practitioners for small- to medium-sized IID tabular data.
Future Work. Our vision for TabArena-v1.0 is a sophisticated benchmark for classification and regression for any tabular dataset for the entire community. Our specific next steps towards this goal are: supporting tiny, large, and non-IID data, integrating new models, and curating more datasets.
To conclude, TabArena is a significant step towards making benchmarking tabular machine learning models a straightforward and reliable process. We look forward to seeing our living benchmark grow and evolve in cooperation with the tabular machine learning community.
Acknowledgments and Disclosure of Funding
We thank the authors of all models with whom we were in contact for feedback and code contributions regarding the implementations of their models, namely, not among the authors of this manuscript: Yury Gorishniy; Paul Koch; Jingang Qu; and Han-Jia Ye. Furthermore, we thank the dataset donors for sharing their data, especially those under a public license. Moreover, we thank the community, particularly OpenML, Kaggle, and Hugging Face, for developing tools that facilitate the sharing of tabular data.
L.P. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under SFB 1597 (SmallData), grant number 499552394. A.T. acknowledges funding by the German Federal Ministry of Education and Research (BMBF). F.H. acknowledges the financial support of the Hector Foundation. This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 539134284, through EFRE (FEIH_2698644) and the state of Baden-Württemberg.
[Uncaptioned image][Uncaptioned image]
Author Contributions
N.E. led the code development, co-led the model curation, led the CPU experiments, contributed to writing the paper, contributed to the management of the collaboration, led the evaluation, and co-led visualizations. L.P. contributed to the code development, contributed to the model curation, led the GPU experiments, led the writing of the paper, led the management of the collaboration, co-led the dataset curation, contributed to the evaluation, and contributed to the visualizations. A.T. helped with the code development, co-led the dataset curation, contributed to writing the paper, contributed to the evaluation, and contributed to the visualizations. D.H. contributed to the code development, co-led the model curation, contributed to writing the paper, contributed to the evaluation, and co-led visualizations. P.M.D. contributed to the code development. D.S. helped with visualizations, helped with supervising the project, and contributed to reviewing and editing the paper. F.H. supervised L.P. and helped with reviewing and editing the paper.
Competing Interests
D.H. is one of the authors of RealMLP and one of the authors of TabICL. D.S. and N.E. are the authors of TabRepo. N.E., L.P., and P.M.D. are developers of AutoGluon, and in extension, the current maintainers of FastAI MLP and Torch MLP. L.P. and F.H. are a subset of the authors of TabPFNv2. L.P. is an OpenML core contributor. F.H. is affiliated with PriorLabs, a company focused on developing tabular foundation models. The authors declare no other competing interests.
Appendix A Additional Experiments
A.1 Alternative Leaderboard Versions
Aggregation Methods.
Here, we provide more leaderboard variants, using different aggregation strategies. Specifically, we obtain errors $\operatorname{err}_{i}$ for each dataset $i$ by averaging error metrics (1-AUROC for binary, logloss for multiclass, and RMSE for regression) over all outer folds. We then aggregate these errors as follows:
  • Elo: As described in Section 2.3.
  • Normalized score: Following [37], we linearly rescale the error such that the best method has a normalized score of one, and the median method has a normalized score of 0. Scores below zero are clipped to zero. These scores are then averaged across datasets.
  • Average rank: Ranks of methods are computed on each dataset (lower is better) and averaged.
  • Harmonic mean rank: Taking the harmonic mean of ranks,
    $\frac{1}{\frac{1}{N}\sum_{i=1}^{N}(1/\text{rank}_{i})},$
    more strongly favors methods having very low ranks on some datasets. It therefore favors methods that are sometimes very good and sometimes very bad over methods that are always mediocre, as the former are more likely to be useful in conjunction with other methods.
  • Improvability: We introduce improvability as a metric that measures how many percent lower the error of the best method is than the current method on a dataset. This is then averaged over datasets. Formally, for a single dataset,
    $\operatorname{Improvability}:=\frac{\operatorname{err}_{i}-\operatorname{best\_err}_{i}}{\operatorname{err}_{i}}\cdot 100\%~{}.$
    Improvability is always between $0\%$ and $100\%$.
Results.
Figure A.1 presents a leaderboard including all models. We impute the results for models on datasets where they are not applicable with the results of RandomForest (default). We choose the default random forest since it is a fast baseline that is sufficiently but not unreasonably weak, to penalize models that are not applicable to all datasets. Table A.1 presents the same data in tabularized format, akin to the current version of the live leaderboard at tabarena.ai. Table A.1 further includes several additional metrics to asses peak average performance, some of which change the ranking (see the color coding) as they are less influenced by model-wise negative outlier results introduced by imputation.
We further investigate our results by presenting the leaderboard across task types. We show the results per task type by computing the results only with datasets from: binary classification in Figure A.2, multiclass classification in Figure A.3, and regression in Figure A.4.
Next, Figure A.5 and Figure A.6 present the results for the TabPFNv2-compatible and TabICL-compatible datasets, but also impute TabPFNv2/TabICL to enable a more direct comparison between these two foundation models. Finally, Figure A.7 presents the result only with datasets for which both TabPFNv2 and TabICL are compatible.
Figure A.1: TabArena-v0.1 leaderboard with imputation for TabPFNv2 and TabICL, Elo (left) and normalized scores (right). For TabPFNv2 and TabICL, on datasets where they are not applicable, we impute their results with RandomForest (default).
Table A.1: TabArena-v0.1 Leaderboard. We show default (D), tuned (T), and tuned + ensembled (T+E) performances. Results of TabPFNv2 and TabICL are imputed with RandomForest (default) for datasets on which they were not run. Times are median times per 1K samples across datasets, averaged over all outer folds per dataset. The best three values in columns are highlighted with gold, silver, and bronze colors. For Elo values, we also indicate their approximate 95% confidence intervals obtained through bootstrapping.
Model Elo ($\uparrow$) Norm. Avg. Harm. #wins ($\uparrow$) Improva- Train time Predict time
score ($\uparrow$) rank ($\downarrow$) mean bility ($\downarrow$) per 1K [s] per 1K [s]
rank ($\downarrow$)
RealMLP (T+E) 1569-27,+26 0.537 8.5 4.9 1 6.4% 6564.71 10.26
TabM (T+E) 1552-29,+21 0.502 9.2 4.4 4 6.7% 3285.87 1.47
LightGBM (T+E) 1532-26,+29 0.452 9.9 5.5 2 8.3% 416.98 2.64
CatBoost (T+E) 1489-23,+25 0.433 11.6 7.3 0 7.6% 1658.41 0.65
CatBoost (T) 1475-27,+22 0.416 12.2 6.1 2 7.8% 1658.41 0.08
TabM (T) 1455-34,+22 0.413 13.0 6.7 1 7.7% 3285.87 0.17
LightGBM (T) 1450-26,+27 0.363 13.1 10.7 0 9.0% 416.98 0.33
XGBoost (T+E) 1446-26,+23 0.361 13.4 8.1 1 9.1% 693.49 1.69
ModernNCA (T+E) 1435-28,+24 0.444 13.9 5.0 3 8.4% 4621.67 8.15
CatBoost (D) 1432-24,+27 0.372 14.0 7.0 2 9.0% 6.83 0.08
TabPFNv2 (T+E) 1414-26,+25 0.476 14.9 3.1 11 8.3% 3030.15 21.44
XGBoost (T) 1409-20,+24 0.313 15.0 12.1 0 9.4% 693.49 0.31
ModernNCA (T) 1404-30,+31 0.307 15.3 8.6 1 9.1% 4621.67 0.47
TabICL (D) 1388-24,+27 0.381 16.0 4.5 6 8.9% 6.63 1.48
RealMLP (T) 1349-21,+26 0.229 17.8 15.1 0 9.8% 6564.71 0.49
TabM (D) 1348-20,+28 0.285 17.8 11.9 0 10.7% 10.49 0.13
TabPFNv2 (T) 1346-23,+22 0.360 18.0 5.6 1 10.4% 3030.15 0.46
TorchMLP (T+E) 1334-18,+29 0.220 18.5 13.4 0 10.3% 2874.67 1.95
TabPFNv2 (D) 1317-27,+28 0.324 19.4 5.5 4 11.4% 3.36 0.31
ModernNCA (D) 1302-23,+23 0.159 20.2 12.2 1 12.6% 14.87 0.31
TabDPT (D) 1298-21,+28 0.300 20.3 4.8 7 12.0% 22.53 8.55
EBM (T+E) 1292-23,+25 0.184 20.6 11.6 1 14.4% 1331.68 0.20
FastaiMLP (T+E) 1250-21,+25 0.158 22.6 13.3 0 13.7% 593.24 4.47
RealMLP (D) 1246-22,+29 0.104 22.9 19.4 0 12.4% 21.86 0.84
ExtraTrees (T+E) 1242-28,+27 0.124 23.0 15.0 0 14.1% 183.02 0.76
EBM (T) 1236-25,+25 0.135 23.3 16.2 0 15.0% 1331.68 0.02
XGBoost (D) 1229-23,+29 0.115 23.5 18.3 0 12.4% 1.94 0.12
TorchMLP (T) 1223-28,+25 0.113 23.8 20.2 0 12.4% 2874.67 0.13
LightGBM (D) 1206-28,+28 0.088 24.8 22.0 0 13.3% 1.96 0.14
EBM (D) 1202-30,+24 0.134 24.9 13.2 1 16.0% 4.67 0.04
RandomForest (T+E) 1201-23,+23 0.098 25.0 12.9 1 14.9% 373.18 0.77
ExtraTrees (T) 1198-28,+22 0.094 25.1 16.6 0 15.1% 183.02 0.09
FastaiMLP (T) 1159-20,+22 0.073 26.8 21.5 0 15.4% 593.24 0.31
RandomForest (T) 1150-26,+25 0.072 27.2 16.6 0 15.9% 373.18 0.09
TorchMLP (D) 1067-26,+24 0.022 30.7 28.2 0 17.2% 9.99 0.13
FastaiMLP (D) 1009-24,+31 0.023 32.7 29.9 0 20.6% 2.86 0.37
RandomForest (D) 1000-0,+0 0.009 33.1 31.3 0 21.0% 0.43 0.05
ExtraTrees (D) 970-23,+36 0.025 34.0 30.3 0 22.8% 0.25 0.05
Linear (T+E) 919-28,+29 0.031 35.6 27.6 0 30.6% 47.49 0.17
Linear (T) 882-28,+23 0.020 36.6 31.9 0 31.4% 47.49 0.07
Linear (D) 856-29,+36 0.014 37.1 29.9 0 32.7% 1.52 0.09
KNN (T+E) 684-27,+23 0.000 40.5 40.2 0 45.3% 2.74 0.18
KNN (T) 607-40,+33 0.000 41.4 41.3 0 46.8% 2.74 0.04
KNN (D) 456-46,+47 0.000 42.8 42.6 0 54.1% 0.05 0.02
Figure A.2: Benchmark results on binary classification with Elo (left) and normalized scores (right). For TabPFNv2 and TabICL, on datasets where they are not applicable, we impute their results with RandomForest (default).
Figure A.3: Benchmark results on multiclass classification with Elo (left) and normalized scores (right). For TabPFNv2 and TabICL, on datasets where they are not applicable, we impute their results with RandomForest (default).
Figure A.4: Benchmark results on regression with Elo (left) and normalized scores (right). For TabPFNv2 and TabICL, on datasets where they are not applicable, we impute their results with RandomForest (default).
Figure A.5: Benchmark results on TabPFNv2-compatible datasets with imputed results for TabICL, using Elo (left) and normalized scores (right). On datasets where TabICL is not applicable, we impute its results with RandomForest (default).
Figure A.6: Benchmark results on TabICL-compatible datasets with imputed results for TabPFNv2, using Elo (left) and normalized scores (right). On datasets where TabPFNv2 is not applicable, we impute its results with RandomForest (default).
Figure A.7: Benchmark results on TabPFNv2- and TabICL-compatible datasets using Elo (left) and normalized scores (right).
A.2 Analyzing Training Time Limit
In our experiments, we restrict the time to evaluate one configuration on one train split of a dataset to $1$ hour. Thus, a model must finish training (across all $8$ inner folds) within $1$ hour, or its training will be gracefully stopped early. Figure A.8 presents the training runtime for all hyperparameter configurations for all models by visualizing what proportion of configurations (x-axis) took how many seconds for training (y-axis).
We observe that for all models, less than $1\%$ of all configurations reach the time limit of $1$ hour. We further investigate the time limit for the GPU-optimized models in Section A.3. For EBMs, we notice that the training was not stopped early at the $1$ hour time limit, positively influencing its results. As this only concerns a small fraction of hyperparameter trials, we did not rerun the training for EBM.
Figure A.8: Training Runtime Analysis, Runtime Distribution (left) and Cumulative Total Runtime (right) Across Hyperparameter Configurations. We show the training runtime in seconds for the hyperparameter configurations across models.
A.3 Tabular Deep Learning on GPU vs. CPU
In our main experiments, we ran TabM and ModernNCA on GPU. To further investigate the impact of hardware choice on these models, we also ran TabM and ModernNCA on CPU. Moreover, we ran RealMLP on GPU. Figure A.9 demonstrates that the hyperparameter configurations of TabM and ModernNCA train much faster and RealMLP slightly faster on GPU than on CPU. We conclude from this ablation that training TabM and ModernNCA on CPU with a time limit of $1$ hour would negatively influence their predictive performance. While the influence is marginal for RealMLP, it seems non-marginal for TabM and ModernNCA.
Figure A.9: Runtime Analysis of RealMLP, TabM, and ModernNCA on CPU vs. GPU. We show the training runtime distribution of RealMLP, TabM, and ModernNCA trained on CPU and GPU across all TabArena datasets. For ModernNCA and TabM on CPU, approximately $16\%$ of runs are early stopped due to the $1$-hour time limit. For GPU, less than $0.1\%$ of runs are early stopped due to the time limit.
A.4 TabArena-Lite
Benchmarking can quickly become very expensive, especially with a sophisticated protocol to guarantee robust results. To reduce the cost of benchmarking, we introduce TabArena-Lite. TabArena-Lite is a continually maintained subset of TabArena that consists, in its first version, of all datasets with one outer fold. Figure A.10 shows results on TabArena-Lite, using 200 hyperparameter configurations per model, but only a single outer fold for all datasets. The results are similar to the results on TabArena in Figure 1, showing that TabArena-Lite is a good indicator of model performance.
To further reduce the cost of benchmarking, we also recommend running new models on TabArena-Lite with one default hyperparameter configuration and optionally with a lower number of random hyperparameter configurations (e.g., $25$). As all other models in TabArena are tuned, a less heavily tuned model that performs comparably could already show that a new model is promising.
We designate TabArena-Lite to be used in academic studies and find any novel model that outperforms other models on at least one dataset, even if it is not among the best on average, a valuable publication. Furthermore, we as maintainers use the performance on TabArena-Lite to prioritize the integration of new models into TabArena. We envision TabArena-Lite also as a living, continuously updated subset. Ideally, future work could determine a method that finds the optimal and most representative subset of partitions and datasets in TabArena to populate TabArena-Lite.
Figure A.10: Benchmark results on TabArena-Lite using Elo (left) and normalized scores (right). Our main leaderboard with TabArena-Lite, a subset of TabArena consisting of all datasets, but only with one outer fold.
A.5 Investigating Statistical Significance
We investigate the statistical significance between models by using critical difference diagrams (CDDs) to represent the results of a Friedman test and then a Nemenyi post-hoc test ($\alpha=0.05$) from AutoRank [^1] https://github.com/sherbold/autorank. Figures A.11, A.12 and A.13 show the CDDs for the full benchmark, TabPFNv2-compatible datasets, and TabICL-compatible datasets with respect to the peak performance of the models, i.e., tuned + ensembled where available. We further investigate statistical significance per-dataset in Appendix E.
We observe that there always exists a group of not statistically significantly different top models containing at least one deep learning model and GBDT, and when available, TabPFNv2 and TabICL.
Figure A.11: Critical Difference Diagram for tuned+ensembled methods on the full benchmark. Lower ranks are better; horizontal bars connect methods that are not statistically significantly different.
Figure A.12: Critical Difference Diagram for tuned+ensembled methods on TabPFNv2-compatible datasets. Lower ranks are better; horizontal bars connect methods that are not statistically significantly different.
Figure A.13: Critical Difference Diagram for tuned+ensembled methods on TabICL-compatible datasets. Lower ranks are better; horizontal bars connect methods that are not statistically significantly different.
Appendix B Data curation
For initializing the TabArena benchmark, we surveyed all the datasets used in 13 previous benchmarking studies: $450$ from PMLB(Mini) [43, 44, 45], $72$ from OpenML-CC18 [29], $45$ from Grinsztajn et al. [32], $11$ from Shwartz-Ziv and Armon [31], $11$ from Gorishniy et al. [30], $176$ from TabZilla [33], $35$ from OpenML-CTR23 [34], $104$ from AMLB [35], $8$ from TabRed [10], $10$ from Tschalzev [8], $279$ from TabRepo [37], $118$ from PyTabKit [20], and $300$ from TALENT [47, 36].
These studies were selected with the goal of covering a wide range of datasets used in tabular benchmarking so that we can clean up the field from problematic or unsuitable datasets. Therefore, each of the studies represents a frequently used benchmark or a general milestone study in the field of tabular machine learning. Combining the dataset collections results in 1053 uniquely named datasets.
For TabArena-v0.1, we aimed at using only those datasets representing realistic, predictive tabular data tasks practitioners would be interested to solve. Therefore, we define a set of selection criteria described in Section B.1. Two of the coauthors manually investigated each of the datasets and applied our selection criteria. We publicly share their notes and curated metadata: tabarena.ai/dataset-curation. Furthermore, we share insights from our curation process in Section B.3
Importantly, we did not exhaustively test each dataset for each of our curation criteria, but proceeded with the next dataset whenever a dataset clearly met at least one of our criteria for exclusion. Therefore, Figure 2 represents the first reason for exclusion that we noticed in a dataset.
Surprisingly, only 51 datasets satisfying all criteria remained. Section B.2 provides additional information on the selected datasets. We consider our data curation a clean-up for tabular data benchmarking that is necessary, but imperfect. Therefore, we aim to continuously improve the data selection and invite researchers to challenge our documented decisions. Section D.3 details protocols to contribute new datasets to TabArena by applying our criteria.
B.1 Dataset Selection Criteria
The datasets for Tabarena-v0.1 were curated by applying the following criteria:
Unique datasets:
We want the TabArena benchmark to be representative of a wide range of tasks without overrepresenting particular tasks. Therefore, we conduct a four-stage deduplication procedure: (1) Automatically filter data sets by name if they match after transforming to the lower case and removing filling characters ’ ’| ’_’ | ’-’. (2) Manually remove datasets where different names were used for the same data set in different studies. (3) Manually remove alternative versions of the same dataset, i.e., temporal data sampled at different rates, or dataset versions with alternative targets. (4) Remove different datasets representing the same task from the same source (i.e., a collection of ML for software tasks named kc1-3).
IID Tabular Data:
We exclude datasets that are non-IID. More specifically, we exclude datasets whose tasks require a non-random split, such as a temporal or group-based split. We leave a non-IID realization of TabArena with temporal and time-series data for future work.
Tabular Domain Tasks:
We exclude datasets from non-tabular modalities transformed into a tabular format. Thus, we exclude featureized image, text, audio, or time series forecasting data. Likewise, we exclude problems that would no longer be solved with tabular machine learning, such as tabular data of control problems solved nowadays by reinforcement learning. While some tasks from other modalities may still be solved using feature extraction and tabular learning methods, it is impossible to assess that without domain experts. Instead of making uninformed decisions, we exclude all datasets from other domains for TabArena-v0.1. In future versions, we consider adding datasets from other domains if there is evidence that tabular learning methods are still a reasonable solution for the task. Therefore, we actively invite researchers from other domains to share datasets for which they apply tabular learning methods.
Real Random Distribution:
We exclude purely artificial data, or any subset thereof, generated by a deterministic function, by sampling from a seeded random process, or by simulating a random distribution. We note that such datasets are still interesting toy functions that help analyze the theoretical capabilities of models qualitatively. Yet, they do not represent distributions from real-world predictive machine learning tasks. While some simulated datasets (i.e., higgs, or MiniBooNE) were conceptualized as machine learning tasks, we decided to exclude them for TabArena-v0.1 for consistency.
Predictive Machine Learning Task:
We exclude tabular data that does not originally stem from a predictive machine learning task for classification or regression. Thus, we exclude tabular data intended for scientific discovery tasks such as anomaly detection, subgroup discovery, data visualization, or causal inference. In particular, this includes survey data never intended for use in a predictive machine learning task. While data from scientific discovery applications can be used for predictive machine learning tasks, we only include it if the original data source intended its use for predictive machine learning, or if a follow-up work re-used the data in a real-world application.
Moreover, we exclude non-predictive tables, where the target label is not predictable based on statistical information from other columns, such as those commonly found in collections like WikiTables [52] or GitTables [53].
We exclude datasets that are trivial to solve and therefore do not represent challenging ML tasks allowing to investigate model differences. We define trivial datasets as datasets where one of the following criteria applies: (1) at least one of the models in our scope is consistently able to achieve perfect performance; (2) multiple models achieve exactly the same highest performance. Note that after applying our set of criteria, none of the considered datasets was found to be trivial.
Size Limit:
We exclude datasets that are tiny or large because they tend to require fundamentally different methods. Tiny datasets require a methodological focus on avoiding overfitting, while methods for large datasets must be very efficient during training. We aim to include tiny and large datasets with dedicated evaluation protocols in future versions of TabArena. For TabArena-v0.1, we exclude datasets with fewer than $500$ or more than $250,000$ samples, measured as the number of training samples after applying our train-test splits. Note that after applying our whole set of criteria, none of the datasets was excluded solely due to being too large, while many datasets were excluded due to a small sample size.
Data Quality:
We exclude datasets that suffer from one of the following data quality issues: (1) heavily preprocessed datasets, such as those where the whole dataset was already used for preprocessing in a way that leaks the target (i.e., PCA); (2) datasets for which we could not find sufficient information to judge their source and preprocessing; (3) datasets with an irreversible target leak. In general, we try to find the original state of the dataset and include it, if applicable. We do not generally exclude preprocessed datasets, as datasets are rarely published without any preprocessing, e.g., due to anonymization. We leave a benchmark with model-specific, domain-specific pre-processing per dataset for future work.
No License Issues:
We exclude any dataset whose license does not allow sharing or using it for an academic benchmark. By doing so, we guard the future of TabArena as a living benchmark, its maintainers, and, most importantly, its users from legal threat.
As a result, we exclude several promising datasets, e.g., due to the default license of Kaggle competitions. Thus, progress towards a less data-restrictive license on Kaggle could greatly benefit the academic community. Likewise, any progress towards sharing more public domain datasets for tabular predictive machine learning would be highly beneficial.
Open-access Structured Data API:
We exclude datasets that cannot be made to be automatically downloaded from a tabular data repository. Eligible data repositories must be open-access, i.e., users do not need an account to download data. Furthermore, the repositories require a structured data and task representation, including metadata information such as feature types, the target column, and outer splits. To the best of our knowledge, only OpenML fulfills these requirements so far. If applicable due to licensing, we manually upload datasets to OpenML to include them in TabArena. This criterion is necessary to enable automated benchmarking and a straightforward user experience.
Ethically Unambiguous Tasks:
We exclude datasets with tasks that pose ethical concerns, such as the Boston Housing dataset [^2] https://fairlearn.org/main/user_guide/datasets/boston_housing_data.html. While curating our datasets, we flagged such datasets and excluded them. We implore the community to investigate our curated datasets for ethical concerns further and immediately notify the maintainers of TabArena about potential problems.
B.2 Included Datasets Details
Table B.1 presents a detailed overview for all datasets included in TabArena-v0.1. We further share all tasks and datasets as an OpenML suite (ID $457$, alias "tabarena-v0.1"). We share BibTex and LaTeX code to easily reference datasets from TabArena in our data curation repository: tabarena.ai/data-tabular-ml-iid-study.
Table B.1: Datasets included in TabArena-v0.1. ’Dataset (Task) ID’ represents the OpenML dataset and task IDs, ’name’ the dataset name, and Ref. the reference corresponding to the dataset. ’N’ represents the no. of samples, ’d’ the no. of features, ’C’ the no. of classes (- for regression tasks), and ’cat’ represents the percentage of features that are categorical. ’Subset’ indicates whether the dataset has been used for the sub-benchmarks focusing on TabPFNv2 (left) and TabICL (right).
Dataset (Task) ID Name Ref. N d C % cat Subset
46913 (363621) blood-transfusion-service-center [54] 748 5 2 20.0 |
46921 (363629) diabetes [55] 768 9 2 11.11 |
46906 (363614) anneal [56] 898 39 5 84.62 |
46954 (363698) QSAR_fish_toxicity [57] 907 7 - 0.0 |
46918 (363626) credit-g [58] 1000 21 2 66.67 |
46941 (363685) maternal_health_risk [59] 1014 7 3 14.29 |
46917 (363625) concrete_compressive_strength [60] 1030 9 - 0.0 |
46952 (363696) qsar-biodeg [61] 1054 42 2 14.29 |
46931 (363675) healthcare_insurance_expenses [62] 1338 7 - 42.86 |
46963 (363707) website_phishing [63] 1353 10 3 100.0 |
46927 (363671) Fitness_Club [64] 1500 7 2 57.14 |
46904 (363612) airfoil_self_noise [65] 1503 6 - 16.67 |
46907 (363615) Another-Dataset-on-used-Fiat-500 [66] 1538 8 - 12.5 |
46980 (363711) MIC [67] 1699 112 8 84.82 |
46938 (363682) Is-this-a-good-customer [68] 1723 14 2 64.29 |
46940 (363684) Marketing_Campaign [69] 2240 26 2 34.62 |
46930 (363674) hazelnut-spread-contaminant-detection [70] 2400 31 2 3.23 |
46956 (363700) seismic-bumps [71] 2584 16 2 25.0 |
46958 (363702) splice [72] 3190 61 3 100.0 |
46912 (363620) Bioresponse [73] 3751 1777 2 0.06 |
46933 (363677) hiva_agnostic [74] 3845 1618 3 100.0 |
46960 (363704) students_dropout_and_academic_success [75] 4424 37 3 48.65 |
46915 (363623) churn [76] 5000 20 2 25.0 |
46953 (363697) QSAR-TID-11 [77] 5742 1025 - 0.0 |
46950 (363694) polish_companies_bankruptcy [78] 5910 65 2 1.54 |
46964 (363708) wine_quality [79] 6497 13 - 7.69 |
46962 (363706) taiwanese_bankruptcy_prediction [80] 6819 95 2 1.05 |
46969 (363689) NATICUSdroid [81] 7491 87 2 100.0 |
46916 (363624) coil2000_insurance_policies [82] 9822 86 2 4.65 |
46911 (363619) Bank_Customer_Churn [83] 10000 11 2 45.45 |
46932 (363676) heloc [84] 10459 24 2 4.17 |
46979 (363712) jm1 [85] 10885 22 2 4.55 |
46924 (363632) E-CommereShippingData [86] 10999 11 2 45.45 |
46947 (363691) online_shoppers_intention [87] 12330 18 2 44.44 |
46937 (363681) in_vehicle_coupon_recommendation [88] 12684 25 2 88.0 |
46942 (363686) miami_housing [89] 13776 16 - 6.25 |
46935 (363679) HR_Analytics_Job_Change_ of_Data_Scientists [90] 19158 13 2 76.92 |
46934 (363678) houses [91] 20640 9 - 0.0 |
46961 (363705) superconductivity [92] 21263 82 - 0.0 |
46919 (363627) credit_card_clients_default [93] 30000 24 2 16.67 |
46905 (363613) Amazon_employee_access [94] 32769 10 2 100.0 |
46910 (363618) bank-marketing [95, 96] 45211 14 2 57.14 |
46928 (363672) Food_Delivery_Time [97] 45451 10 - 30.0 |
46949 (363693) physiochemical_protein [98] 45730 10 - 0.0 |
46939 (363683) kddcup09_appetency [99] 50000 213 2 18.31 |
46923 (363631) diamonds [100] 53940 10 - 30.0 |
46922 (363630) Diabetes130US [101] 71518 48 2 83.33 |
46908 (363616) APSFailure [102] 76000 171 2 0.58 |
46955 (363699) SDSS17 [103] 78053 12 3 25.0 |
46920 (363628) customer_satisfaction_in_airline [104] 129880 22 2 77.27 |
46929 (363673) GiveMeSomeCredit [105] 150000 11 2 9.09 |
B.3 Noteworthy Observations from Curation
We observed several trends while curating the datasets for TabArena-v0.1. To improve the discussion related to datasets in our community, we share some noteworthy trends below.
  • For various datasets, it was not possible to automate the selection process, because the metadata that would be required is not available. Therefore, given the current state of data repositories, we consider that automated curation procedures produce more biased results than careful manual curation. Finding out which splits are appropriate for a task, or whether the targets were created using deterministic functions, requires substantial effort and oftentimes, reading and understanding the papers where datasets were introduced. To still make the inclusion of datasets as objective as possible, we introduce a checklist for new datasets in Section D.3.
  • Most of the datasets excluded due to license issues were Kaggle datasets with restrictive licenses, which otherwise would have been well-suited for inclusion. In the future, we hope that more high-quality datasets with less restrictive licenses will become available, also on Kaggle.
  • The large amount of datasets from other modalities seems to be an artifact from times before the development of high-performing modality-specific approaches. At least 16 datasets were images for handwritten digit or letter recognition. As those tasks are clearly outdated, we excluded them. To be consistent, we also excluded datasets consisting of features from image data for which we were not able to assess whether the tasks are outdated. Features extracted from image data are not an exclusion criterion for datasets in TabArena, as long as they represent a meaningful task and tabular models are a reasonable approach to solve those tasks.
  • The huge amount of tiny datasets is likely an artifact of a time when data collection was done at a much smaller scale than nowadays. Only four of the 142 tiny datasets for which we found a publication date were published later than 2010. Moreover, many of the tiny datasets seem to have originated in educational content, such as books or toy examples in tutorials.
  • Several datasets used in previous benchmarking studies were originally introduced as part of a series of AutoML Challenges. Datasets in these challenges were often released (and shared) with obscured, non-meaningful names. Most of the datasets are ablated versions of other datasets, and therefore have led to unintended duplicates in existing benchmarks. Furthermore, many of those datasets were from other domains, like images or text.
  • Of the 254 datasets with alternative versions listed in Figure 2, most are from the PMLB benchmark [44] and represent differently parameterized versions of artificially created datasets: 118 are Feynman equations and 62 are Friedman data generation functions.
  • Throughout the benchmarks, inconsistent versions of the same datasets were used: tasks were binarized, features were removed, and sometimes even targets were changed. This can be partially attributed to the misleading versioning system of OpenML. Subsequent versions of the same datasets correspond to a different upload with the same name, not necessarily an improved version of the same datasets. Therefore, some studies reused the alternative versions of the dataset uploaded under the same name for specific studies. In gathering the datasets, we disregarded which version of a dataset was used and solely focused on names. Therefore, some alternative versions were already filtered for the set of 1053 datasets with unique names. In our benchmark, we always searched for the raw version and used the dataset with minimal preprocessing.
  • After applying all other criteria, only 51 datasets were found to satisfy the IID criterion, while 68 did not. This underscores the findings of [10] that all previous benchmarks used random splits inappropriately. TabArena aims to end this malpractice.
  • A large number of datasets are tabular but were not intended to be used for predictive tasks. Most of these datasets were filtered due to being ’scientific discovery’ tasks, some due to quality issues. In general, some of these datasets might still be useful for benchmarking if they represent realistic distributions and target functions. However, most of the datasets filtered due to this criterion appeared to be relatively simple tasks. That is, some were already found to be trivially solvable in other studies, and some contained only a few features. In the future, we are open to considering including datasets not initially intended for predictive tasks, if no other issues are found, and if one can argue for potential predictive machine learning applications.
Appendix C Model Curation
C.1 Implementation Framework Details
We implement models (and their unit tests) based on the AbstractModel framework [^3] https://auto.gluon.ai/stable/tutorials/tabular/advanced/tabular-custom-model.html from AutoGluon [19]. In particular, we implement model-specific preprocessing, training, and inference within the AbstractModel framework for all models. The framework allows us to use all functionalities from AutoGluon, TabRepo, and in extension scikit-learn [24] to run models in a standardized way. Moreover, the pipeline logic encompassing models within TabArena is implemented in a tested, sophisticated framework that is regularly used in real-world applications.
To integrate models in AbstractModel framework, we require two properties of a model implementation: (I) Iteratively trained models (e.g., GBDTs or MLPs) must support early stopping based on a time limit. Moreover, they must support the use of externally provided validation data. (II) We require a default model-specific preprocessing pipeline that handles, if needed, data anomalies such as NaN values, categorical features, or feature scaling. The model-agnostic preprocessing of the AbstractModel framework detects categorical features, transforms text or image features, and cleans common data problems.
The original implementations of some models do not fulfill these requirements; thus, we added support ourselves or together with the model authors. Our requirements aim to get the most out of models in a proper benchmark and in real-world pipelines. Early stopping based on a time limit avoids model failures due to time constraints in benchmarks and is quintessential for integration into time-constrained, real-world pipelines. Likewise, only models that support externally provided validation data can be properly used in pipelines with pre-defined validation splits. Finally, different models require different preprocessing, and relying only on one shared model-agnostic preprocessing pipeline is inappropriate. We detail the model-agnostic and model-specific below.
Foundation Model Implementation Details. For all foundation models, we refit on training and validation data instead of using cross-validation ensembles, following recommendations from the authors of TabPFN and TabICL. We hypothesize that the foundation models do not gain much from cross-validation ensembles because, unlike other models, they do not utilize the validation data per fold for early stopping during training. Thus, their in-context learning might benefit more from using the training and validation data as context for inference on test data.
The foundation models TabPFNv2 and TabICL have been released with restrictions in terms of dataset size. In particular, TabPFNv2 is restricted to datasets with up to $10,000$ training samples, $500$ features, and $10$ classes for classification tasks. TabICL is constrained to classification tasks with up to $100,000$ training samples and $500$ features. TabDPT has no size restrictions because it natively relies on context retrieval, dimensionality reduction, and class codes during inference [23]. For context retrieval, we use the default context size of $1024$ described in the paper. Thereby, we override the implementation’s default of $128$, which we found to perform poorly in preliminary experiments.
We use the newest available checkpoints for all foundation models. For TabDPT, we use tabdpt1_1.pth. For TabICL, we use tabicl-classifier-v1.1-0506.ckpt. For TabPFN, we use the defaults for classification tabpfn-v2-classifier.ckpt and regression tabpfn-v2-regressor.ckpt, as well as all other checkpoints during HPO: tabpfn-v2-classifier-gn2p4bpt.ckpt, tabpfn-v2-classifier-llderlii.ckpt, tabpfn-v2-classifier-od3j1g5m.ckpt, tabpfn-v2-classifier-vutqq28w.ckpt, tabpfn-v2-classifier-znskzxi4.ckpt, tabpfn-v2-regressor-09gpqh39.ckpt, tabpfn-v2-regressor-2noar4o2.ckpt, tabpfn-v2-regressor-5wof9ojf.ckpt, tabpfn-v2-regressor-wyl4o83o.ckpt.
Model-agnostic Preprocessing.
Our model-agnostic preprocessing relies on AutoGluon’s AutoMLPipelineFeatureGenerator [^4] https://auto.gluon.ai/stable/tutorials/tabular/tabular-feature-engineering.html. The model-agnostic preprocessing can handle boolean, numerical, categorical, datetime, and text columns. Importantly, the implementation of a model can control how the model-agnostic preprocessing treats the input data. As a result, a model could obtain raw text and datetime columns as input, such that its model-specific preprocessing can handle them.
For TabArena, we let the model-agnostic preprocessing handle text and datetime columns. Text columns are transformed to n-hot encoded n-grams. Datetime columns are converted into a Pandas datetime and into multiple columns representing the year, month, day, and day of the week. Numerical columns are left untouched. Categorical columns are replaced with categorical codes to save memory space. The columns are, nevertheless, further treated as categorical. Finally, constant or duplicated columns are dropped. Importantly, we always keep missing values and delegate handling them to the model-specific preprocessing.
Model-specific Preprocessing.
We perform minimal invasive model-specific preprocessing and otherwise rely on the preprocessing already implemented within the model’s code. Specifically, we use the following model-specific preprocessing before passing the data to the model’s code:
  • CatBoost, LightGBM, XGBoost, EBM, TabICL, TabPFNv2, FastaiMLP, and TorchMLP do not use any custom model-specific preprocessing and rely entirely on the model’s code.
  • RandomForest and ExtraTrees use ordinal encoding for categorical variables. Missing values are imputed to 00.
  • TabDPT uses ordinal encoding for categorical variables.
  • RealMLP handles missing numericals by mean imputation with a missingness indicator.
  • TabM and ModernNCA use the numerical quantile-based preprocessing from TabM and then use mean imputation with an indicator for numerical features.
  • Linear uses one-hot-encoding, mean or median imputation (hyperparameter), standard scaling, and quartile transformation (hyperparameter).
  • KNN drops all categorical features and fills missing numerical values with 00. Moreover, it uses leave-one-out cross-validation instead of $8$-fold cross-validation. The leave-one-out cross-validation is natively implemented into the KNN model logic and allows for obtaining the validation predictions per sample very efficiently.
C.2 Hyperparameter search spaces
In the following, we list some details and hyperparameter search spaces for different models:
  • The search spaces for CatBoost (Table C.1), LightGBM (Table C.2), XGBoost (Table C.3), RandomForest (Table C.4), and ExtraTrees (Table C.5) were determined based on experiments. We assessed a large set of hyperparameters inspired by the respective documentations as well as several papers [e.g., 37, 106, 20] and experimentally determined good ranges for them for tuning. We verified that the new search spaces outperform the original search spaces on TabRepo [37].
    For gradient-boosted trees, we use the implementation from AutoGluon with its early stopping logic and n_estimators=10_000. Compared to TabRepo, which used 300 overall estimators for Random Forest and ExtraTrees and used out-of-bag predictions for validation, we fit these models using 8-fold CV with 50 estimators per model, resulting in 400 estimators overall.
  • For EBM(Table C.6), we use a search space provided by the authors.
  • For RealMLP (Table C.7), we also use a search space provided by the authors. For the default parameters, we turn off label smoothing since we are not using accuracy as our evaluation metric, as recommended by [20].
  • For TabM (Table C.8) and ModernNCA (Table C.9), we use search spaces coordinated with the authors. For the batch size, we choose a training-set-size dependent logic following the TabM paper [9]:
[…]
  • FastaiMLP (Table C.10) and TorchMLP (Table C.11) were taken from AutoGluon, in dialogue with the maintainers/authors.
  • Linear (Table C.12) and KNN (Table C.13) were taken from TabRepo.
  • For TabPFNv2, we use the search space from the original paper and the official repository, in coordination with the authors.
  • For TabICL and TabDPT, we only use their default configurations. For TabICL and TabDPT, we use the newest checkpoint (see Section C.1), unlike the original paper.
Table C.1: Hyperparameter search space for CatBoost.
Hyperparameter Space
learning_rate LogUniform([0.005, 0.1])
bootstrap_type Bernoulli
subsample Uniform([0.7, 1.0])
grow_policy Choice(["SymmetricTree", "Depthwise"])
depth UniformInt([4, 8])
colsample_bylevel Uniform([0.85, 1.0])
l2_leaf_reg LogUniform([1e-4, 5])
leaf_estimation_iterations LogUniformInt([1, 20])
one_hot_max_size LogUniformInt([8, 100])
model_size_reg LogUniform([0.1, 1.5])
max_ctr_complexity UniformInt([2, 5])
boosting_type Plain
max_bin 254
Table C.2: Hyperparameter search space for LightGBM.
Hyperparameter Space
learning_rate LogUniform([0.005, 0.1])
feature_fraction Uniform([0.4, 1.0])
bagging_fraction Uniform([0.7, 1.0])
bagging_freq 1
num_leaves LogUniformInt([2, 200])
min_data_in_leaf LogUniformInt([1, 64])
extra_trees Choice([False, True])
min_data_per_group LogUniformInt([2, 100])
cat_l2 LogUniform([0.005, 2])
cat_smooth LogUniform([0.001, 100])
max_cat_to_onehot LogUniformInt([8, 100])
lambda_l1 Uniform([1e-4, 1.0])
lambda_l2 Uniform([1e-4, 2.0])
Table C.3: Hyperparameter search space for XGBoost.
Hyperparameter Space
learning_rate LogUniform([0.005, 0.1])
max_depth LogUniformInt([4, 10])
min_child_weight LogUniform([0.001, 5.0])
subsample Uniform([0.6, 1.0])
colsample_bylevel Uniform([0.6, 1.0])
colsample_bynode Uniform([0.6, 1.0])
reg_alpha Uniform([1e-4, 5.0])
reg_lambda Uniform([1e-4, 5.0])
grow_policy Choice(["depthwise", "lossguide"])
max_cat_to_onehot LogUniformInt([8, 100])
max_leaves LogUniformInt([8, 1024])
Table C.4: Hyperparameter search space for Random Forest.
Hyperparameter Space
max_features Uniform([0.4, 1.0])
max_samples Uniform([0.5, 1.0])
min_samples_split LogUniformInt([2, 4])
bootstrap Choice([False, True])
n_estimators 50
min_impurity_decrease LogUniform([1e-5, 1e-3])
Table C.5: Hyperparameter search space for ExtraTrees.
Hyperparameter Space
max_features Choice(["sqrt", 0.5, 0.75, 1.0])
min_samples_split LogUniformInt([2, 32])
bootstrap False
n_estimators 50
min_impurity_decrease Choice([0.0, 1e-5, 3e-5, 1e-4, 3e-4, 1e-3], p=[0.5, 0.1, 0.1, 0.1, 0.1, 0.1])
Table C.6: Hyperparameter search space for EBM.
Hyperparameter Space
max_leaves UniformInt([2, 3])
smoothing_rounds UniformInt([0, 1000])
learning_rate LogUniform([0.0025, 0.2])
interactions Uniform([0.95, 0.999])
interaction_smoothing_rounds UniformInt([0, 200])
min_hessian LogUniform([1e-10, 1e-2])
min_samples_leaf UniformInt([2, 20])
validation_size Uniform([0.05, 0.25])
early_stopping_tolerance LogUniform([1e-10, 1e-5])
gain_scale LogUniform([0.5, 5.0])
Table C.7: Hyperparameter search space for RealMLP. With probability 0.5, either the “Default” or the “Large” option is chosen for each configuration.
Hyperparameter Space
n_hidden_layers UniformInt([2, 4])
hidden_sizes rectangular
hidden_width Choice([256, 384, 512])
p_drop Uniform([0.0, 0.5])
act mish
plr_sigma LogUniform([1e-2, 50])
sq_mom 1 -- LogUniform([0.005, 0.05])
plr_lr_factor LogUniform([0.05, 0.3])
scale_lr_factor LogUniform([2.0, 10.0])
first_layer_lr_factor LogUniform([0.3, 1.5])
ls_eps_sched coslog4
ls_eps LogUniform([0.005, 0.1])
lr LogUniform([0.02, 0.3])
wd LogUniform([0.001, 0.05])
use_ls Choice([False, True])
early_stopping_additive_patience 40
early_stopping_multiplicative_patience 3
Default (prob=0.5) Large (prob=0.5)
plr_hidden_1 16 Choice([8, 16, 32, 64])
plr_hidden_2 4 Choice([8, 16, 32, 64])
n_epochs 256 Choice([256, 512])
use_early_stopping False True
Table C.8: Hyperparameter search space for TabM.
Hyperparameter Space
batch_size auto
patience 16
amp False
arch_type tabm-mini
tabm_k 32
gradient_clipping_norm 1.0
share_training_batches False
lr LogUniform([1e-4, 3e-3])
weight_decay Choice([0.0, LogUniform([1e-4, 1e-1])])
n_blocks UniformInt([2, 5])
d_block Choice([128, 144, 160, …, 1008, 1024])
dropout Choice([0.0, Uniform([0.0, 0.5])])
num_emb_type pwl
d_embedding Choice([8, 12, 16, 20, 24, 28, 32])
num_emb_n_bins UniformInt([2, 128])
Table C.9: Hyperparameter search space for ModernNCA.
Hyperparameter Space
dropout Uniform([0.0, 0.5])
d_block UniformInt([64, 1024])
n_blocks Choice([0, UniformInt([0, 2])])
dim UniformInt([64, 1024])
num_emb_n_frequencies UniformInt([16, 96])
num_emb_frequency_scale LogUniform([0.005, 10.0])
num_emb_d_embedding UniformInt([16, 64])
sample_rate Uniform([0.05, 0.6])
lr LogUniform([1e-5, 1e-1])
weight_decay Choice([0.0, LogUniform([1e-6, 1e-3])])
temperature 1.0
num_emb_type plr
num_emb_lite True
batch_size auto
Table C.10: Hyperparameter search space for FastaiMLP.
Hyperparameter Space
layers Choice([[200], [400], [200, 100], [400, 200], [800, 400], [200, 100, 50], [400, 200, 100]])
emb_drop Uniform([0.0, 0.7])
ps Uniform([0.0, 0.7])
bs Choice([128, 256, 512, 1024, 2048])
lr LogUniform([5e-4, 1e-1])
epochs UniformInt([20, 50])
Table C.11: Hyperparameter search space for TorchMLP.
Hyperparameter Space
learning_rate LogUniform([1e-4, 3e-2])
weight_decay LogUniform([1e-12, 0.1])
dropout_prob Uniform([0.0, 0.4])
use_batchnorm Choice([False, True])
num_layers UniformInt([1, 5])
hidden_size UniformInt([8, 256])
activation Choice(["relu", "elu"])
Table C.12: Hyperparameter search space for LinearModel.
Hyperparameter Space
C Uniform([0.1, 1000])
proc.skew_threshold Choice([0.9, 0.99, 0.999, None])
proc.impute_strategy Choice(["median", "mean"])
penalty Choice(["L2", "L1"])
Table C.13: Hyperparameter search space for KNN.
Hyperparameter Space
n_neighbors Choice([3, 4, 5, 6, 7, 8, 9, 10, 11, 13, 15, 20, 30, 40, 50])
weights Choice(["uniform", "distance"])
p Choice([2, 1])
C.3 TabArena Ensemble
The TabArena ensemble highlighted in Figure 6 was created by ensembling a portfolio, a set of hyperparameter configurations across models. Given a portfolio, we evaluate each of its models in sequence until a time limit is reached or all models have been evaluated. Then, we post-hoc ensemble [1] all evaluated hyperparameter configurations. For the sake of Figure 6, we simulated the TabArena ensemble using the result artifacts.
We created a portfolio following the learning procedure introduced by [37] using leave-one-dataset-out cross-validation with a portfolio of size 200 and 40 ensemble selection steps. We leave further discussion and investigation of portfolio learning with the results of TabArena to future work.
Appendix D Using and Contributing to the Living Benchmark
D.1 Benchmarking with TabArena
To benchmark a model, a user must (1) implement their model in the AbstractModel framework; (2) create a search space; (3) run the implementation on TabArena or TabArena-Lite; (4) and analyze the results. We provide code and more detailed documentation for these three steps in our code repositories with examples: tabarena.ai/code-examples. Below, we provide a snapshot [^5] Parts of this snapshot may become outdated due to the benchmarking system being updated. of code snippets for each step: model implementation (Listing LABEL:lst:impl_model), search space (Listing LABEL:lst:search_space), benchmarking (Listing LABEL:lst:benchmarking), and analysis of the results (Listing LABEL:lst:res).
Listing 1: Implementing a custom RandomForest model for TabArena.
1import numpy as np
2from autogluon.core.models import AbstractModel
3from autogluon.features import LabelEncoderFeatureGenerator
4
5class CustomRandomForestModel(AbstractModel):
6 ag_key = "CRF"
7 ag_name = "CustomRF"
8
9 def __init__(self, **kwargs):
10 super().__init__(**kwargs)
11 self._feature_generator = None
12
13 def _preprocess(self, X: pd.DataFrame, is_train=False, **kwargs) -> np.ndarray:
14 """Model-specific preprocessing of the input data."""
15 X = super()._preprocess(X, **kwargs)
16 if is_train:
17 self._feature_generator = LabelEncoderFeatureGenerator(verbosity=0)
18 self._feature_generator.fit(X=X)
19 if self._feature_generator.features_in:
20 X = X.copy()
21 X[self._feature_generator.features_in] = self._feature_generator.transform(
22 X=X
23 )
24 return X.fillna(0).to_numpy(dtype=np.float32)
25
26 def _fit(self, X, y):
27 from sklearn.ensemble import RandomForestRegressor, RandomForestClassifier
28 if self.problem_type in ["regression"]:
29 model_cls = RandomForestRegressor
30 else:
31 model_cls = RandomForestClassifier
32
33 X = self.preprocess(X, is_train=True)
34 self.model = model_cls(**self._get_model_params())
35 self.model.fit(X, y)
Listing 2: Creating a search space for the custom RandomForest model.
1def get_configs_for_custom_rf(num_random_configs):
2 from autogluon.common.space import Int
3 from tabrepo.utils.config_utils import ConfigGenerator
4
5 gen_custom_rf = ConfigGenerator(
6 model_cls=CustomRandomForestModel,
7 manual_configs=[{}],
8 search_space== {
9 "n_estimators": Int(4, 50),
10 },
11 )
12 return gen_custom_rf.generate_all_bag_experiments(
13 num_random_configs=num_random_configs
14 )
Listing 3: Benchmarking the custom RandomForest model.
1import openml
2from tabrepo.benchmark.experiment import run_experiments_new
3
4task_ids = openml.study.get_suite("tabarena-v0.1").tasks
5methods = get_configs_for_custom_rf(num_random_configs=1)
6
7run_experiments_new(
8 output_dir="/path/to/output/dir",
9 model_experiments=model_experiments,
10 tasks=task_ids,
11 repetitions_mode="TabArena-Lite",
12)
Listing 4: Comparing the custom RandomForest model to the leaderboard.
1import pandas as pd
2from tabrepo.paper.paper_runner_tabarena import PaperRunTabArena
3
4from . import post_process_local_results
5from . import load_local_results, load_paper_reuslts
6
7repo = post_process_local_results()
8plotter = PaperRunTabArena(repo=repo, output_dir=EVAL_DIR)
9
10df_results = load_local_results(plotter)
11df_results = load_paper_reuslts(df_results)
12
13# Create and save the leaderboard figure and table
14plotter.eval(
15 df_results=df_results,
16 framework_types_extra=list(df_results["config_type"].unique()),
17)
D.2 Contributing Models
To include a new model in TabArena, we ask users to open an issue on the TabArena benchmarking code repository (tabarena.ai/code) to start the process of adding a model. We envision this process not as a static request but as an ongoing interaction between the contributors and maintainers. During this process, the goal is to populate the issue over time with the information necessary to integrate a model. We require the following information to include a new model:
  1. 1.
    Public Model Implementation. The model must be implemented in the AbstractModel framework (see Section C.1), the code for this implementation must be publicly shared, and it must pass the default unit test for TabArena models. The implementation should represent a standalone model and not, for example, an ensembling pipeline of several existing models or sub-calls to other machine learning systems. We leave benchmarking for such pipelines, or in general, machine learning systems, to future iterations of TabArena.
    Finally, note that the model can also first be implemented in a scikit-learn API-like interface and then wrapped with the AbstractModel framework. This would be the recommended workflow in many cases.
  2. 2.
    Preprocessing and Hyperparameters. The implementation should specify model-specific preprocessing (see Section C.1). Moreover, the contributor must recommend default hyperparameters and a search space for hyperparameter optimization.
  3. 3.
    Model Verification. The maintainers of TabArena must have reviewed the source code of the model. In an ideal process, this review could also help the user to improve their model and implementation. In addition, the model should demonstrate promising results on TabArena-Lite. Moreover, if the contributor is not among the original authors of the model, the contributor (potentially in coordination with the maintainers of TabArena) shall reach out to the original authors to verify the implementation and its optimal intended usage. This may involve including the original author in GitHub issues, reviewing the pull request, or validating the results.
  4. 4.
    Maintenance Commitment. While the TabArena team generally maintains model implementations, we might need help from the original contributors to resolve future version conflicts or outdated functionality. Therefore, contributors must share their preferred way of being contacted. Note that the TabArena team may deprecate models that are no longer maintainable, consistently outperformed by newer models, or have bugs that cannot be reasonably resolved.
Once the issue is deemed finalized, two maintainers of TabArena need to review and approve the issue to complete the model integration.
D.3 Contributing Data: New Datasets and Curation Feedback
We envision TabArena to be a platform for discussing benchmarking practices. Therefore, we invite users, researchers, and practitioners to challenge our curation decisions or provide curation feedback using GitHub issues in the TabArena curation repository: tabarena.ai/data-tabular-ml-iid-study.
Moreover, we also invite the community to add new datasets. For a new dataset to be added to TabArena, we outline the current template below:
  1. 1.
    Reference to pull request with a .yaml file including a dataset description following the template in the repository.
  2. 2.
    Reference to a .py file containing a preprocessing pipeline to transform data from the raw data source into a format suitable for benchmarking.
  3. 3.
    A checklist answering the following questions
    1. (a)
      Is the data available through an API for automatic downloading, or does the license allow for reuploading the data?
    2. (b)
      What is the sample size?
    3. (c)
      Was the data extracted from another modality (i.e., text, image, time-series)
      1. If yes: Are tabular learning methods a reasonable solution compared to domain-specific methods? (If possible, provide a reference)
    4. (d)
      Is there a deterministic function for optimally mapping the features to the target?
    5. (e)
      Was the data generated artificially or from a parameterized simulation?
    6. (f)
      Can you provide a one-sentence user story detailing the benefits of better predictive performance in this task?
    7. (g)
      Were the samples collected over time?
      1. If yes: Is the task about predicting future data, and, if yes, are there distribution shifts for samples collected later?
    8. (h)
      Were the samples collected in different groups (i.e., transactions from different customers, patients from multiple hospitals, repeated experimental results from different batches)?
      1. If yes: Is the task about predicting samples from unseen groups, and if yes, are distribution shifts of samples from different groups expected?
    9. (i)
      Are there known preprocessing techniques already applied to the ‘rawest’ available data version?
    10. (j)
      What preprocessing steps are recommended to conceptualize the task in the preprocessing Python file?
    11. (k)
      Do you have any other recommendations for how to use the dataset for benchmarking?
The maintainers will verify the provided information and engage in review-like discussions if necessary. After verifying that the task is reasonable, the dataset will be included in the next benchmark version.
The above protocol outlines the user-driven process for adding new datasets. However, we welcome any suggestions for datasets that could be included in future versions of TabArena and where we, as maintainers, drive the process to add the dataset. For that, we welcome GitHub issues with the ‘Dataset Suggestion’ template, which includes: (1) a link to the raw data, and (2) the dataset license. The TabArena maintainers will review the suggested dataset by applying the outlined protocol and, if the criteria are met, include it in the next version of TabArena.
The checklist results from our learnings during data curation and covers the essential aspects where we had to look closely at the data in our curation process. However, we want to emphasize that we do not generally exclude datasets using this checklist. On the contrary, for future versions of TabArena, we aim to explicitly extend the benchmark with tasks that are not covered sufficiently so far, either due to a lack of high-quality data or due to a lack of domain knowledge to judge the task quality on our end. Therefore, we encourage users to propose datasets from other domains, non-IID data, and for any supervised learning task consisting of tabular features where strong performance is a desired property.
D.3.1 Checklist Examples
In the following, we provide examples of the application of our checklist to one included and one excluded dataset.
Example for the APSFailure dataset, which represents one of the borderline cases that were included in TabArena-v0.1:
  1. a)
    Is the data available through an API for automatic downloading, or does the license allow for reuploading the data?    Yes
  2. b)
    What is the sample size?    76,000
  3. c)
    Was the data extracted from another modality (i.e., text, image, time-series)?    Unclear, as the data was anonymized. Some features represent histograms, so some of the features possibly were extracted from time-series.
    1. If yes: Are tabular learning methods a reasonable solution compared to domain-specific methods? (If possible, provide reference)    The data is from a 2016 challenge and was provided by a well-known company. Given that the dataset is comparably recent and the source is legitimate, we conclude that it still represents a meaningful tabular data task.
  4. d)
    Is there a deterministic function for optimally mapping the features to the target?    No
  5. e)
    Was the data generated artificially or from a parameterized simulation?    No
  6. f)
    Can you provide a one-sentence user story detailing the benefits of a better predictive performance in this task?    By automatically detecting component failures in trucks, the company can save costly manual effort and prevent accidents from releasing trucks with faulty components.
  7. g)
    Were the samples collected over time? Probably yes.
    1. If yes: Is the task about predicting future data, and, if yes, are there distribution shifts for samples collected later? In a real application, future data would be predicted, however, the provided test dataset revealed that no distribution shifts between train and test data can be expected as the features are time-invariant.
  8. h)
    Were the samples collected in different groups (i.e. transactions from different customers, patients from multiple hospitals, repeated experimental results from different batches)?    No
    1. If yes: Is the task about predicting samples from unseen groups, and if yes, are distribution shifts of samples from different groups expected?    N/A
  9. i)
    Are there known preprocessing techniques that have already been applied to the ‘rawest’ available data version?    The feature names were anonymized. Some features were preprocessed.
  10. j)
    What preprocessing steps are recommended to conceptualize the task in the preprocessing Python file?    Combine the original training and test files. Convert "na" strings to real NaN/missing values for numeric features.
  11. k)
    Do you have any other recommendations for how to use the dataset for benchmarking?    The data originally comes with a cost-matrix, which could be considered in future benchmark versions.
Example for the socmob dataset, which was excluded for TabArena-v0.1 as it represents a scientific discovery task where higher predictive performance is not relevant:
  1. a)
    Is the data available through an API for automatic downloading, or does the license allow for reuploading the data?    Yes
  2. b)
    What is the sample size?    1156
  3. c)
    Was the data extracted from another modality (i.e., text, image, time-series)    No
    1. If yes: Are tabular learning methods a reasonable solution compared to domain-specific methods? (If possible, provide reference)    N/A
  4. d)
    Is there a deterministic function for optimally mapping the features to the target?    No
  5. e)
    Was the data generated artificially or from a parameterized simulation?    No
  6. f)
    Can you provide a one-sentence user story detailing the benefits of a better predictive performance in this task?    No. The data was collected to empirically test the hypothesis that associations between socioeconomic and occupational attributes of fathers and sons among sons from intact families are stronger than associations between attributes of fathers and sons among sons from any kind of disrupted or reconstituted families. The dataset has one target and five predictive features, including the investigated family structure. Although supervised (linear) models are applied to the data, the goal is not to maximize performance, but to empirically quantify the relationship of one feature to the target while controlling for confounding factors (other features).
  7. g)
    Were the samples collected over time?    No, the study was cross-sectional and collected data in 1973.
    1. If yes: Is the task about predicting future data, and, if yes, are there distribution shifts for samples collected later?    N/A
  8. h)
    Were the samples collected in different groups (i.e. transactions from different customers, patients from multiple hospitals, repeated experimental results from different batches)?    No
    1. If yes: Is the task about predicting samples from unseen groups, and if yes, are distribution shifts of samples from different groups expected?    N/A
  9. i)
    Are there known preprocessing techniques that have already been applied to the ‘rawest’ available data version?    No noteworthy steps.
  10. j)
    What preprocessing steps are recommended to conceptualize the task in the preprocessing Python file?    None.
  11. k)
    Do you have any other recommendations for how to use the dataset for benchmarking?    Do not use the data for benchmarking the capabilities of predictive modeling approaches, but maybe for a scientific discovery benchmark in the future.
D.4 Contributing Results: Leaderboard Submissions
We seek to define a process for TabArena to submit to the leaderboard that satisfies the following principles: (1) Equality: Submitting to the leaderboard is accessible in the same way to everyone. (2) Transparency: All attempts to submit to the leaderboard are transparent to the public. (3) Reproducibility: Submitted results are reproducible. (4) Fairness: Cheated results, i.e., by utilizing the test data in an inappropriate way or simply by submitting manually altered results, are rejected. (5) : Feasibility: The submission process, in particular the validation, must be manageable for the maintainers in a reasonable amount of time.
Using these guiding principles, we define our submission process:
  1. 1.
    To submit results to the leaderboard, users can write a pull request to tabarena.ai/community-results that contains:
    1. (a)
      An update to the results dataset collection with new data for their model.
    2. (b)
      Reproducible and documented code to obtain the results. We require users to start the process to add their new model to TabArena (as described in Section D.2) and to train and evaluate their approach using the provided TabArena benchmarking code.
    3. (c)
      A description or link to a description, e.g., a paper, for the new model.
    4. (d)
      The following statement: "I confirm that these results were produced using the attached modeling pipeline and to the best of my knowledge, I have used the test data appropriately and have not manipulated the results."
    5. (e)
      Indicate whether verification of the submitted results by the maintainers of TabArena is requested.
  2. 2.
    The maintainers will verify that all the required information is present and will proceed depending on whether verification was requested:
    1. (a)
      Non-verified submission (fast): The request will be merged without recomputing the results. Non-verified submissions will not appear on the landing page and will be presented as a separate leaderboard on tabarena.ai [^6] Note that for TabArena-v0.1 no non-validated leaderboard exists on the website. This will change with the first submission from the community using this protocol..
    2. (b)
      Verified submission: The maintainers will manually review the code and reproduce the results for a random sample of outer folds from different datasets. If the results can be reproduced successfully and no further issues are found, the request will be merged and the results will appear in the main TabArena leaderboard.
We aim to continuously improve our submission process and welcome any feedback or suggestions for future versions of TabArena.
D.5 Running TabArena Models in Practice
Models integrated into TabArena can be easily used to solve predictive machine learning tasks on new datasets, independent of the TabArena benchmark. Listing LABEL:lst:run_model shows how to run RealMLP on a toy dataset from scikit-learn. For more details on this code, please see our code repositories with examples: tabarena.ai/code-examples.
Listing 5: Running RealMLP from TabArena on a new dataset.
1from autogluon.core.data import LabelCleaner
2from autogluon.features.generators import AutoMLPipelineFeatureGenerator
3from sklearn.datasets import load_breast_cancer
4from sklearn.metrics import roc_auc_score
5from sklearn.model_selection import train_test_split
6
7# Import a TabArena model
8from tabrepo.benchmark.models.ag.realmlp.realmlp_model import RealMLPModel
9
10# Get Data
11X, y = load_breast_cancer(return_X_y=True, as_frame=True)
12X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.5, random_state=42)
13
14# Model-agnostic Preprocessing
15feature_generator, label_cleaner = AutoMLPipelineFeatureGenerator(), LabelCleaner.construct(problem_type="binary", y=y)
16X_train, y_train = feature_generator.fit_transform(X_train), label_cleaner.transform(y_train)
17X_test, y_test = feature_generator.transform(X_test), label_cleaner.transform(y_test)
18
19# Train TabArena Model
20clf = RealMLPModel()
21clf.fit(X=X_train, y=y_train)
22
23# Predict and score
24prediction_probabilities = clf.predict_proba(X=X_test)
25print("ROC AUC:", roc_auc_score(y_test, prediction_probabilities))
Appendix E Performance Results Per Dataset
Table E.1 presents the performance per dataset for all methods in TabArena-v0.1.
Table E.1: Performance Per Dataset.  We show the average predictive performance per dataset with the standard deviation over folds.  We show the performance for the default hyperparameter configuration (Default), for the model after tuning (Tuned), and for the ensemble after tuning (Tuned + Ens.).  We highlight the best-performing methods with significance on three levels:  (1) Green: The best performing method on average;  (2) Bold: Methods that are not significantly worse than the best method on average, based on a Wilcoxon Signed-Rank test for paired samples with Holm-Bonferroni correction and $\alpha=0.05$.  (3) Underlined: Methods that are not significantly worse than the best method in the same pipeline regime (Default, Tuned, or Tuned + Ens.), based on a Wilcoxon Signed-Rank test for paired samples with Holm-Bonferroni correction and $\alpha=0.05$. We exclude AutoGluon for significance tests in the Tuned + Ens. regime.
APSFailure (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.990 $\pm$ 0.002 0.990 $\pm$ 0.002 0.990 $\pm$ 0.002
ExtraTrees 0.990 $\pm$ 0.002 0.990 $\pm$ 0.003 0.991 $\pm$ 0.002
XGBoost 0.992 $\pm$ 0.002 0.992 $\pm$ 0.002 0.992 $\pm$ 0.002
LightGBM 0.992 $\pm$ 0.002 0.992 $\pm$ 0.002 0.992 $\pm$ 0.002
CatBoost 0.992 $\pm$ 0.003 0.992 $\pm$ 0.002 0.992 $\pm$ 0.003
EBM 0.991 $\pm$ 0.002 0.991 $\pm$ 0.002 0.991 $\pm$ 0.002
FastAIMLP 0.988 $\pm$ 0.003 0.989 $\pm$ 0.002 0.991 $\pm$ 0.002
TorchMLP 0.990 $\pm$ 0.002 0.991 $\pm$ 0.002 0.992 $\pm$ 0.001
RealMLP 0.991 $\pm$ 0.002 0.991 $\pm$ 0.002 0.992 $\pm$ 0.002
TabM 0.992 $\pm$ 0.002 0.992 $\pm$ 0.002 0.993 $\pm$ 0.002
MNCA 0.991 $\pm$ 0.003 0.991 $\pm$ 0.002 0.993 $\pm$ 0.002
TabPFNv2 - - -
TabDPT 0.990 $\pm$ 0.003 - -
TabICL 0.993 $\pm$ 0.002 - -
Linear 0.988 $\pm$ 0.002 0.987 $\pm$ 0.003 0.989 $\pm$ 0.001
KNN 0.910 $\pm$ 0.011 0.975 $\pm$ 0.004 0.979 $\pm$ 0.004
AutoGluon - - 0.993 $\pm$ 0.002
Amazon_employee_access (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.839 $\pm$ 0.005 0.841 $\pm$ 0.005 0.849 $\pm$ 0.005
ExtraTrees 0.833 $\pm$ 0.006 0.841 $\pm$ 0.007 0.845 $\pm$ 0.006
XGBoost 0.834 $\pm$ 0.007 0.859 $\pm$ 0.008 0.862 $\pm$ 0.008
LightGBM 0.843 $\pm$ 0.009 0.850 $\pm$ 0.007 0.858 $\pm$ 0.009
CatBoost 0.882 $\pm$ 0.008 0.883 $\pm$ 0.008 0.883 $\pm$ 0.007
EBM 0.839 $\pm$ 0.006 0.841 $\pm$ 0.007 0.842 $\pm$ 0.006
FastAIMLP 0.854 $\pm$ 0.007 0.853 $\pm$ 0.008 0.866 $\pm$ 0.007
TorchMLP 0.835 $\pm$ 0.007 0.838 $\pm$ 0.006 0.849 $\pm$ 0.007
RealMLP 0.844 $\pm$ 0.007 0.846 $\pm$ 0.008 0.864 $\pm$ 0.008
TabM 0.833 $\pm$ 0.009 0.842 $\pm$ 0.009 0.849 $\pm$ 0.009
MNCA 0.846 $\pm$ 0.008 0.860 $\pm$ 0.008 0.869 $\pm$ 0.007
TabPFNv2 - - -
TabDPT 0.841 $\pm$ 0.006 - -
TabICL 0.854 $\pm$ 0.006 - -
Linear 0.848 $\pm$ 0.009 0.848 $\pm$ 0.009 0.851 $\pm$ 0.008
KNN 0.839 $\pm$ 0.005 0.839 $\pm$ 0.005 0.839 $\pm$ 0.005
AutoGluon - - 0.882 $\pm$ 0.005
Another-Dataset-on-used-Fiat-500 (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 750.8 $\pm$ 28.4 736.5 $\pm$ 24.8 735.4 $\pm$ 25.3
ExtraTrees 744.2 $\pm$ 29.5 735.8 $\pm$ 26.5 735.1 $\pm$ 26.7
XGBoost 754.6 $\pm$ 23.0 741.4 $\pm$ 22.2 737.1 $\pm$ 22.6
LightGBM 746.0 $\pm$ 22.4 740.4 $\pm$ 24.6 729.4 $\pm$ 22.7
CatBoost 738.1 $\pm$ 20.8 736.3 $\pm$ 21.8 732.9 $\pm$ 21.9
EBM 749.9 $\pm$ 22.9 750.0 $\pm$ 24.1 745.7 $\pm$ 23.6
FastAIMLP 760.5 $\pm$ 17.6 761.2 $\pm$ 21.6 756.3 $\pm$ 21.4
TorchMLP 775.0 $\pm$ 26.3 769.8 $\pm$ 25.0 765.0 $\pm$ 24.7
RealMLP 757.4 $\pm$ 23.8 756.0 $\pm$ 22.4 726.6 $\pm$ 24.0
TabM 752.1 $\pm$ 23.3 755.0 $\pm$ 25.0 748.0 $\pm$ 22.8
MNCA 753.7 $\pm$ 27.0 748.2 $\pm$ 27.6 731.1 $\pm$ 28.2
TabPFNv2 727.7 $\pm$ 23.8 733.2 $\pm$ 27.2 727.4 $\pm$ 26.0
TabDPT 724.0 $\pm$ 22.1 - -
TabICL - - -
Linear 793.8 $\pm$ 25.4 764.7 $\pm$ 19.5 764.9 $\pm$ 19.6
KNN 882.0 $\pm$ 23.7 862.4 $\pm$ 30.5 852.8 $\pm$ 28.9
AutoGluon - - 729.6 $\pm$ 24.6
Bank_Customer_Churn (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.851 $\pm$ 0.008 0.857 $\pm$ 0.008 0.858 $\pm$ 0.008
ExtraTrees 0.851 $\pm$ 0.006 0.860 $\pm$ 0.007 0.861 $\pm$ 0.007
XGBoost 0.864 $\pm$ 0.009 0.866 $\pm$ 0.010 0.866 $\pm$ 0.010
LightGBM 0.864 $\pm$ 0.009 0.866 $\pm$ 0.009 0.867 $\pm$ 0.009
CatBoost 0.870 $\pm$ 0.009 0.870 $\pm$ 0.009 0.870 $\pm$ 0.009
EBM 0.862 $\pm$ 0.010 0.864 $\pm$ 0.010 0.864 $\pm$ 0.010
FastAIMLP 0.859 $\pm$ 0.009 0.863 $\pm$ 0.007 0.864 $\pm$ 0.008
TorchMLP 0.860 $\pm$ 0.008 0.866 $\pm$ 0.008 0.866 $\pm$ 0.008
RealMLP 0.866 $\pm$ 0.009 0.870 $\pm$ 0.009 0.871 $\pm$ 0.009
TabM 0.869 $\pm$ 0.010 0.871 $\pm$ 0.010 0.871 $\pm$ 0.010
MNCA 0.864 $\pm$ 0.008 0.869 $\pm$ 0.008 0.869 $\pm$ 0.008
TabPFNv2 0.872 $\pm$ 0.009 0.874 $\pm$ 0.009 0.874 $\pm$ 0.008
TabDPT 0.859 $\pm$ 0.008 - -
TabICL 0.868 $\pm$ 0.009 - -
Linear 0.772 $\pm$ 0.010 0.773 $\pm$ 0.010 0.773 $\pm$ 0.010
KNN 0.526 $\pm$ 0.005 0.558 $\pm$ 0.009 0.557 $\pm$ 0.008
AutoGluon - - 0.869 $\pm$ 0.009
Bioresponse (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.873 $\pm$ 0.007 0.873 $\pm$ 0.007 0.876 $\pm$ 0.006
ExtraTrees 0.867 $\pm$ 0.008 0.868 $\pm$ 0.009 0.871 $\pm$ 0.008
XGBoost 0.873 $\pm$ 0.008 0.875 $\pm$ 0.008 0.876 $\pm$ 0.008
LightGBM 0.872 $\pm$ 0.008 0.874 $\pm$ 0.007 0.875 $\pm$ 0.008
CatBoost 0.872 $\pm$ 0.009 0.875 $\pm$ 0.008 0.874 $\pm$ 0.008
EBM 0.852 $\pm$ 0.009 0.863 $\pm$ 0.008 0.866 $\pm$ 0.008
FastAIMLP 0.850 $\pm$ 0.011 0.857 $\pm$ 0.010 0.860 $\pm$ 0.010
TorchMLP 0.846 $\pm$ 0.008 0.856 $\pm$ 0.009 0.863 $\pm$ 0.008
RealMLP 0.858 $\pm$ 0.009 0.864 $\pm$ 0.008 0.875 $\pm$ 0.008
TabM 0.863 $\pm$ 0.005 0.871 $\pm$ 0.007 0.873 $\pm$ 0.007
MNCA 0.860 $\pm$ 0.010 0.865 $\pm$ 0.007 0.874 $\pm$ 0.008
TabPFNv2 - - -
TabDPT 0.862 $\pm$ 0.010 - -
TabICL - - -
Linear 0.789 $\pm$ 0.011 0.789 $\pm$ 0.011 0.793 $\pm$ 0.013
KNN 0.805 $\pm$ 0.009 0.840 $\pm$ 0.010 0.849 $\pm$ 0.009
AutoGluon - - 0.878 $\pm$ 0.007
Diabetes130US (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.631 $\pm$ 0.006 0.656 $\pm$ 0.008 0.657 $\pm$ 0.008
ExtraTrees 0.623 $\pm$ 0.009 0.651 $\pm$ 0.007 0.653 $\pm$ 0.008
XGBoost 0.662 $\pm$ 0.008 0.668 $\pm$ 0.008 0.670 $\pm$ 0.008
LightGBM 0.648 $\pm$ 0.008 0.668 $\pm$ 0.007 0.672 $\pm$ 0.008
CatBoost 0.671 $\pm$ 0.008 0.671 $\pm$ 0.008 0.672 $\pm$ 0.008
EBM 0.659 $\pm$ 0.008 0.662 $\pm$ 0.007 0.665 $\pm$ 0.008
FastAIMLP 0.647 $\pm$ 0.007 0.656 $\pm$ 0.008 0.662 $\pm$ 0.008
TorchMLP 0.655 $\pm$ 0.007 0.663 $\pm$ 0.009 0.667 $\pm$ 0.009
RealMLP 0.659 $\pm$ 0.005 0.662 $\pm$ 0.008 0.669 $\pm$ 0.008
TabM 0.660 $\pm$ 0.007 0.662 $\pm$ 0.008 0.663 $\pm$ 0.008
MNCA 0.658 $\pm$ 0.008 0.662 $\pm$ 0.008 0.666 $\pm$ 0.008
TabPFNv2 - - -
TabDPT 0.609 $\pm$ 0.008 - -
TabICL 0.647 $\pm$ 0.008 - -
Linear 0.648 $\pm$ 0.008 0.648 $\pm$ 0.008 0.648 $\pm$ 0.008
KNN 0.516 $\pm$ 0.004 0.539 $\pm$ 0.005 0.541 $\pm$ 0.005
AutoGluon - - 0.673 $\pm$ 0.008
E-CommereShippingData (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.739 $\pm$ 0.005 0.740 $\pm$ 0.006 0.741 $\pm$ 0.005
ExtraTrees 0.737 $\pm$ 0.006 0.741 $\pm$ 0.006 0.740 $\pm$ 0.005
XGBoost 0.740 $\pm$ 0.006 0.742 $\pm$ 0.007 0.742 $\pm$ 0.006
LightGBM 0.739 $\pm$ 0.006 0.740 $\pm$ 0.005 0.741 $\pm$ 0.006
CatBoost 0.744 $\pm$ 0.006 0.742 $\pm$ 0.007 0.741 $\pm$ 0.007
EBM 0.744 $\pm$ 0.004 0.743 $\pm$ 0.005 0.743 $\pm$ 0.004
FastAIMLP 0.737 $\pm$ 0.005 0.741 $\pm$ 0.007 0.741 $\pm$ 0.007
TorchMLP 0.737 $\pm$ 0.008 0.740 $\pm$ 0.007 0.741 $\pm$ 0.007
RealMLP 0.741 $\pm$ 0.007 0.742 $\pm$ 0.007 0.742 $\pm$ 0.007
TabM 0.744 $\pm$ 0.008 0.740 $\pm$ 0.005 0.743 $\pm$ 0.006
MNCA 0.741 $\pm$ 0.010 0.741 $\pm$ 0.006 0.741 $\pm$ 0.004
TabPFNv2 0.744 $\pm$ 0.007 0.744 $\pm$ 0.007 0.744 $\pm$ 0.006
TabDPT 0.735 $\pm$ 0.007 - -
TabICL 0.743 $\pm$ 0.006 - -
Linear 0.704 $\pm$ 0.006 0.721 $\pm$ 0.008 0.722 $\pm$ 0.008
KNN 0.731 $\pm$ 0.005 0.737 $\pm$ 0.007 0.738 $\pm$ 0.007
AutoGluon - - 0.738 $\pm$ 0.005
Fitness_Club (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.775 $\pm$ 0.018 0.801 $\pm$ 0.015 0.800 $\pm$ 0.015
ExtraTrees 0.769 $\pm$ 0.018 0.816 $\pm$ 0.013 0.815 $\pm$ 0.014
XGBoost 0.798 $\pm$ 0.015 0.808 $\pm$ 0.015 0.808 $\pm$ 0.015
LightGBM 0.795 $\pm$ 0.015 0.815 $\pm$ 0.015 0.814 $\pm$ 0.015
CatBoost 0.814 $\pm$ 0.014 0.812 $\pm$ 0.015 0.811 $\pm$ 0.015
EBM 0.813 $\pm$ 0.015 0.810 $\pm$ 0.015 0.812 $\pm$ 0.015
FastAIMLP 0.806 $\pm$ 0.013 0.814 $\pm$ 0.013 0.814 $\pm$ 0.014
TorchMLP 0.813 $\pm$ 0.016 0.813 $\pm$ 0.017 0.814 $\pm$ 0.016
RealMLP 0.812 $\pm$ 0.015 0.814 $\pm$ 0.015 0.816 $\pm$ 0.014
TabM 0.818 $\pm$ 0.014 0.818 $\pm$ 0.014 0.818 $\pm$ 0.014
MNCA 0.817 $\pm$ 0.014 0.814 $\pm$ 0.015 0.801 $\pm$ 0.017
TabPFNv2 0.822 $\pm$ 0.012 0.820 $\pm$ 0.013 0.817 $\pm$ 0.013
TabDPT 0.818 $\pm$ 0.014 - -
TabICL 0.819 $\pm$ 0.013 - -
Linear 0.819 $\pm$ 0.014 0.818 $\pm$ 0.015 0.818 $\pm$ 0.015
KNN 0.733 $\pm$ 0.019 0.802 $\pm$ 0.014 0.803 $\pm$ 0.014
AutoGluon - - 0.811 $\pm$ 0.012
Food_Delivery_Time (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 7.855 $\pm$ 0.041 7.587 $\pm$ 0.046 7.588 $\pm$ 0.046
ExtraTrees 8.179 $\pm$ 0.045 7.753 $\pm$ 0.053 7.749 $\pm$ 0.053
XGBoost 7.397 $\pm$ 0.055 7.397 $\pm$ 0.055 7.400 $\pm$ 0.055
LightGBM 7.616 $\pm$ 0.053 7.378 $\pm$ 0.054 7.374 $\pm$ 0.053
CatBoost 7.379 $\pm$ 0.051 7.367 $\pm$ 0.051 7.368 $\pm$ 0.051
EBM 7.433 $\pm$ 0.040 7.424 $\pm$ 0.051 7.412 $\pm$ 0.044
FastAIMLP 8.188 $\pm$ 0.053 8.085 $\pm$ 0.056 8.060 $\pm$ 0.052
TorchMLP 7.735 $\pm$ 0.059 7.579 $\pm$ 0.049 7.559 $\pm$ 0.052
RealMLP 7.926 $\pm$ 0.053 7.453 $\pm$ 0.051 7.414 $\pm$ 0.046
TabM 7.760 $\pm$ 0.050 7.651 $\pm$ 0.053 7.651 $\pm$ 0.052
MNCA 7.486 $\pm$ 0.044 7.421 $\pm$ 0.045 7.379 $\pm$ 0.047
TabPFNv2 - - -
TabDPT 7.551 $\pm$ 0.042 - -
TabICL - - -
Linear 8.563 $\pm$ 0.047 8.325 $\pm$ 0.065 8.271 $\pm$ 0.055
KNN 8.552 $\pm$ 0.047 8.256 $\pm$ 0.042 8.162 $\pm$ 0.049
AutoGluon - - 7.362 $\pm$ 0.051
GiveMeSomeCredit (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.846 $\pm$ 0.003 0.862 $\pm$ 0.003 0.863 $\pm$ 0.003
ExtraTrees 0.840 $\pm$ 0.003 0.857 $\pm$ 0.002 0.857 $\pm$ 0.003
XGBoost 0.865 $\pm$ 0.002 0.866 $\pm$ 0.002 0.866 $\pm$ 0.002
LightGBM 0.865 $\pm$ 0.002 0.866 $\pm$ 0.002 0.867 $\pm$ 0.002
CatBoost 0.866 $\pm$ 0.002 0.867 $\pm$ 0.002 0.867 $\pm$ 0.002
EBM 0.864 $\pm$ 0.002 0.865 $\pm$ 0.002 0.865 $\pm$ 0.002
FastAIMLP 0.829 $\pm$ 0.004 0.843 $\pm$ 0.005 0.847 $\pm$ 0.003
TorchMLP 0.863 $\pm$ 0.002 0.864 $\pm$ 0.002 0.865 $\pm$ 0.002
RealMLP 0.865 $\pm$ 0.002 0.866 $\pm$ 0.002 0.866 $\pm$ 0.002
TabM 0.866 $\pm$ 0.002 0.867 $\pm$ 0.002 0.867 $\pm$ 0.002
MNCA 0.866 $\pm$ 0.002 0.866 $\pm$ 0.002 0.867 $\pm$ 0.002
TabPFNv2 - - -
TabDPT 0.842 $\pm$ 0.003 - -
TabICL 0.866 $\pm$ 0.002 - -
Linear 0.841 $\pm$ 0.002 0.841 $\pm$ 0.002 0.841 $\pm$ 0.002
KNN 0.570 $\pm$ 0.003 0.672 $\pm$ 0.003 0.673 $\pm$ 0.003
AutoGluon - - 0.867 $\pm$ 0.002
HR_Analytics_Job_Change (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.789 $\pm$ 0.006 0.802 $\pm$ 0.006 0.802 $\pm$ 0.006
ExtraTrees 0.784 $\pm$ 0.007 0.800 $\pm$ 0.007 0.801 $\pm$ 0.007
XGBoost 0.805 $\pm$ 0.006 0.803 $\pm$ 0.005 0.805 $\pm$ 0.006
LightGBM 0.802 $\pm$ 0.007 0.803 $\pm$ 0.007 0.804 $\pm$ 0.006
CatBoost 0.804 $\pm$ 0.006 0.804 $\pm$ 0.006 0.804 $\pm$ 0.006
EBM 0.800 $\pm$ 0.006 0.800 $\pm$ 0.006 0.801 $\pm$ 0.006
FastAIMLP 0.801 $\pm$ 0.005 0.801 $\pm$ 0.007 0.803 $\pm$ 0.007
TorchMLP 0.801 $\pm$ 0.006 0.801 $\pm$ 0.006 0.803 $\pm$ 0.006
RealMLP 0.801 $\pm$ 0.007 0.801 $\pm$ 0.006 0.803 $\pm$ 0.006
TabM 0.801 $\pm$ 0.007 0.803 $\pm$ 0.006 0.803 $\pm$ 0.006
MNCA 0.801 $\pm$ 0.006 0.802 $\pm$ 0.007 0.803 $\pm$ 0.007
TabPFNv2 - - -
TabDPT 0.801 $\pm$ 0.006 - -
TabICL 0.805 $\pm$ 0.006 - -
Linear 0.796 $\pm$ 0.006 0.796 $\pm$ 0.006 0.796 $\pm$ 0.006
KNN 0.605 $\pm$ 0.009 0.663 $\pm$ 0.003 0.672 $\pm$ 0.003
AutoGluon - - 0.805 $\pm$ 0.007
Is-this-a-good-customer (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.721 $\pm$ 0.020 0.727 $\pm$ 0.018 0.729 $\pm$ 0.020
ExtraTrees 0.695 $\pm$ 0.021 0.713 $\pm$ 0.022 0.718 $\pm$ 0.024
XGBoost 0.723 $\pm$ 0.021 0.742 $\pm$ 0.023 0.744 $\pm$ 0.022
LightGBM 0.724 $\pm$ 0.020 0.741 $\pm$ 0.022 0.746 $\pm$ 0.020
CatBoost 0.748 $\pm$ 0.020 0.743 $\pm$ 0.019 0.744 $\pm$ 0.019
EBM 0.751 $\pm$ 0.019 0.745 $\pm$ 0.018 0.748 $\pm$ 0.018
FastAIMLP 0.711 $\pm$ 0.018 0.742 $\pm$ 0.025 0.745 $\pm$ 0.017
TorchMLP 0.728 $\pm$ 0.020 0.727 $\pm$ 0.023 0.733 $\pm$ 0.018
RealMLP 0.732 $\pm$ 0.023 0.731 $\pm$ 0.025 0.742 $\pm$ 0.020
TabM 0.744 $\pm$ 0.022 0.743 $\pm$ 0.019 0.744 $\pm$ 0.019
MNCA 0.738 $\pm$ 0.024 0.732 $\pm$ 0.020 0.705 $\pm$ 0.024
TabPFNv2 0.746 $\pm$ 0.019 0.735 $\pm$ 0.022 0.743 $\pm$ 0.018
TabDPT 0.742 $\pm$ 0.016 - -
TabICL 0.744 $\pm$ 0.019 - -
Linear 0.738 $\pm$ 0.021 0.737 $\pm$ 0.024 0.736 $\pm$ 0.023
KNN 0.498 $\pm$ 0.026 0.534 $\pm$ 0.034 0.534 $\pm$ 0.028
AutoGluon - - 0.745 $\pm$ 0.019
MIC (logloss $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.513 $\pm$ 0.031 0.485 $\pm$ 0.023 0.474 $\pm$ 0.019
ExtraTrees 0.521 $\pm$ 0.030 0.482 $\pm$ 0.020 0.470 $\pm$ 0.018
XGBoost 0.470 $\pm$ 0.020 0.440 $\pm$ 0.019 0.440 $\pm$ 0.019
LightGBM 0.503 $\pm$ 0.019 0.453 $\pm$ 0.020 0.453 $\pm$ 0.019
CatBoost 0.455 $\pm$ 0.020 0.453 $\pm$ 0.019 0.451 $\pm$ 0.018
EBM 0.475 $\pm$ 0.018 0.446 $\pm$ 0.016 0.445 $\pm$ 0.016
FastAIMLP 0.506 $\pm$ 0.024 0.462 $\pm$ 0.023 0.450 $\pm$ 0.020
TorchMLP 0.473 $\pm$ 0.019 0.465 $\pm$ 0.024 0.453 $\pm$ 0.017
RealMLP 0.492 $\pm$ 0.027 0.439 $\pm$ 0.021 0.434 $\pm$ 0.017
TabM 0.432 $\pm$ 0.017 0.431 $\pm$ 0.017 0.430 $\pm$ 0.016
MNCA 0.465 $\pm$ 0.019 0.455 $\pm$ 0.020 0.451 $\pm$ 0.018
TabPFNv2 0.468 $\pm$ 0.043 0.440 $\pm$ 0.022 0.433 $\pm$ 0.022
TabDPT 0.481 $\pm$ 0.021 - -
TabICL 0.465 $\pm$ 0.022 - -
Linear 0.589 $\pm$ 0.035 0.589 $\pm$ 0.035 0.589 $\pm$ 0.035
KNN 2.040 $\pm$ 0.075 1.105 $\pm$ 0.096 1.019 $\pm$ 0.096
AutoGluon - - 0.445 $\pm$ 0.018
Marketing_Campaign (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.883 $\pm$ 0.015 0.881 $\pm$ 0.016 0.882 $\pm$ 0.015
ExtraTrees 0.884 $\pm$ 0.015 0.886 $\pm$ 0.015 0.888 $\pm$ 0.015
XGBoost 0.897 $\pm$ 0.015 0.903 $\pm$ 0.016 0.904 $\pm$ 0.015
LightGBM 0.901 $\pm$ 0.014 0.911 $\pm$ 0.015 0.911 $\pm$ 0.014
CatBoost 0.907 $\pm$ 0.015 0.904 $\pm$ 0.014 0.903 $\pm$ 0.015
EBM 0.903 $\pm$ 0.015 0.905 $\pm$ 0.015 0.906 $\pm$ 0.016
FastAIMLP 0.890 $\pm$ 0.017 0.905 $\pm$ 0.015 0.909 $\pm$ 0.014
TorchMLP 0.898 $\pm$ 0.015 0.910 $\pm$ 0.014 0.915 $\pm$ 0.013
RealMLP 0.906 $\pm$ 0.015 0.907 $\pm$ 0.014 0.911 $\pm$ 0.014
TabM 0.901 $\pm$ 0.016 0.916 $\pm$ 0.013 0.916 $\pm$ 0.014
MNCA 0.909 $\pm$ 0.016 0.912 $\pm$ 0.016 0.909 $\pm$ 0.015
TabPFNv2 0.915 $\pm$ 0.015 0.915 $\pm$ 0.013 0.919 $\pm$ 0.013
TabDPT 0.896 $\pm$ 0.016 - -
TabICL 0.911 $\pm$ 0.013 - -
Linear 0.906 $\pm$ 0.013 0.906 $\pm$ 0.013 0.905 $\pm$ 0.013
KNN 0.591 $\pm$ 0.021 0.615 $\pm$ 0.022 0.628 $\pm$ 0.020
AutoGluon - - 0.915 $\pm$ 0.013
NATICUSdroid (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.977 $\pm$ 0.002 0.981 $\pm$ 0.003 0.981 $\pm$ 0.003
ExtraTrees 0.977 $\pm$ 0.002 0.982 $\pm$ 0.002 0.982 $\pm$ 0.002
XGBoost 0.985 $\pm$ 0.002 0.985 $\pm$ 0.002 0.985 $\pm$ 0.002
LightGBM 0.985 $\pm$ 0.002 0.986 $\pm$ 0.002 0.986 $\pm$ 0.002
CatBoost 0.986 $\pm$ 0.002 0.986 $\pm$ 0.002 0.986 $\pm$ 0.002
EBM 0.984 $\pm$ 0.002 0.985 $\pm$ 0.001 0.985 $\pm$ 0.001
FastAIMLP 0.985 $\pm$ 0.002 0.985 $\pm$ 0.001 0.986 $\pm$ 0.001
TorchMLP 0.985 $\pm$ 0.002 0.985 $\pm$ 0.001 0.986 $\pm$ 0.002
RealMLP 0.985 $\pm$ 0.002 0.986 $\pm$ 0.001 0.986 $\pm$ 0.001
TabM 0.986 $\pm$ 0.001 0.986 $\pm$ 0.001 0.986 $\pm$ 0.001
MNCA 0.983 $\pm$ 0.002 0.984 $\pm$ 0.002 0.985 $\pm$ 0.002
TabPFNv2 0.983 $\pm$ 0.002 0.984 $\pm$ 0.003 0.985 $\pm$ 0.002
TabDPT 0.985 $\pm$ 0.002 - -
TabICL 0.987 $\pm$ 0.001 - -
Linear 0.981 $\pm$ 0.002 0.981 $\pm$ 0.002 0.981 $\pm$ 0.002
KNN 0.977 $\pm$ 0.002 0.977 $\pm$ 0.002 0.977 $\pm$ 0.002
AutoGluon - - 0.987 $\pm$ 0.002
QSAR-TID-11 (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.806 $\pm$ 0.047 0.805 $\pm$ 0.048 0.796 $\pm$ 0.046
ExtraTrees 0.806 $\pm$ 0.046 0.802 $\pm$ 0.046 0.791 $\pm$ 0.046
XGBoost 0.786 $\pm$ 0.049 0.761 $\pm$ 0.047 0.760 $\pm$ 0.048
LightGBM 0.772 $\pm$ 0.050 0.758 $\pm$ 0.049 0.756 $\pm$ 0.048
CatBoost 0.774 $\pm$ 0.049 0.773 $\pm$ 0.048 0.771 $\pm$ 0.049
EBM 0.872 $\pm$ 0.039 0.859 $\pm$ 0.042 0.853 $\pm$ 0.042
FastAIMLP 0.776 $\pm$ 0.045 0.766 $\pm$ 0.049 0.761 $\pm$ 0.050
TorchMLP 0.774 $\pm$ 0.052 0.762 $\pm$ 0.049 0.748 $\pm$ 0.054
RealMLP 0.763 $\pm$ 0.047 0.764 $\pm$ 0.049 0.754 $\pm$ 0.050
TabM 0.761 $\pm$ 0.050 0.758 $\pm$ 0.050 0.755 $\pm$ 0.049
MNCA 0.771 $\pm$ 0.045 0.744 $\pm$ 0.044 0.734 $\pm$ 0.043
TabPFNv2 - - -
TabDPT 0.773 $\pm$ 0.046 - -
TabICL - - -
Linear 1.020 $\pm$ 0.032 0.940 $\pm$ 0.040 0.937 $\pm$ 0.039
KNN 0.806 $\pm$ 0.047 0.806 $\pm$ 0.047 0.806 $\pm$ 0.047
AutoGluon - - 0.747 $\pm$ 0.049
QSAR_fish_toxicity (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.907 $\pm$ 0.047 0.885 $\pm$ 0.050 0.884 $\pm$ 0.049
ExtraTrees 0.880 $\pm$ 0.052 0.873 $\pm$ 0.055 0.870 $\pm$ 0.052
XGBoost 0.905 $\pm$ 0.050 0.881 $\pm$ 0.043 0.879 $\pm$ 0.042
LightGBM 0.894 $\pm$ 0.043 0.889 $\pm$ 0.045 0.883 $\pm$ 0.044
CatBoost 0.877 $\pm$ 0.049 0.875 $\pm$ 0.045 0.874 $\pm$ 0.047
EBM 0.905 $\pm$ 0.050 0.904 $\pm$ 0.048 0.898 $\pm$ 0.047
FastAIMLP 0.908 $\pm$ 0.048 0.909 $\pm$ 0.046 0.897 $\pm$ 0.048
TorchMLP 0.906 $\pm$ 0.055 0.897 $\pm$ 0.055 0.890 $\pm$ 0.055
RealMLP 0.878 $\pm$ 0.054 0.884 $\pm$ 0.058 0.865 $\pm$ 0.052
TabM 0.910 $\pm$ 0.047 0.898 $\pm$ 0.051 0.887 $\pm$ 0.048
MNCA 0.882 $\pm$ 0.053 0.882 $\pm$ 0.053 0.873 $\pm$ 0.053
TabPFNv2 0.868 $\pm$ 0.047 0.873 $\pm$ 0.051 0.860 $\pm$ 0.049
TabDPT 0.859 $\pm$ 0.049 - -
TabICL - - -
Linear 0.950 $\pm$ 0.056 0.950 $\pm$ 0.056 0.950 $\pm$ 0.056
KNN 0.962 $\pm$ 0.064 0.935 $\pm$ 0.061 0.919 $\pm$ 0.065
AutoGluon - - 0.880 $\pm$ 0.054
SDSS17 (logloss $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.085 $\pm$ 0.002 0.073 $\pm$ 0.002 0.072 $\pm$ 0.002
ExtraTrees 0.131 $\pm$ 0.002 0.080 $\pm$ 0.002 0.078 $\pm$ 0.002
XGBoost 0.074 $\pm$ 0.002 0.074 $\pm$ 0.002 0.074 $\pm$ 0.002
LightGBM 0.087 $\pm$ 0.002 0.073 $\pm$ 0.003 0.073 $\pm$ 0.002
CatBoost 0.075 $\pm$ 0.002 0.074 $\pm$ 0.003 0.074 $\pm$ 0.003
EBM 0.087 $\pm$ 0.002 0.081 $\pm$ 0.002 0.081 $\pm$ 0.002
FastAIMLP 0.134 $\pm$ 0.004 0.111 $\pm$ 0.004 0.112 $\pm$ 0.003
TorchMLP 0.094 $\pm$ 0.003 0.081 $\pm$ 0.002 0.080 $\pm$ 0.002
RealMLP 0.103 $\pm$ 0.002 0.089 $\pm$ 0.002 0.087 $\pm$ 0.002
TabM 0.097 $\pm$ 0.002 0.083 $\pm$ 0.002 0.083 $\pm$ 0.002
MNCA 0.082 $\pm$ 0.002 0.076 $\pm$ 0.002 0.075 $\pm$ 0.002
TabPFNv2 - - -
TabDPT 0.088 $\pm$ 0.001 - -
TabICL 0.076 $\pm$ 0.002 - -
Linear 0.146 $\pm$ 0.003 0.147 $\pm$ 0.003 0.139 $\pm$ 0.003
KNN 1.155 $\pm$ 0.030 0.512 $\pm$ 0.006 0.440 $\pm$ 0.006
AutoGluon - - 0.067 $\pm$ 0.002
airfoil_self_noise (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 1.898 $\pm$ 0.095 1.891 $\pm$ 0.101 1.851 $\pm$ 0.096
ExtraTrees 1.822 $\pm$ 0.094 1.676 $\pm$ 0.106 1.683 $\pm$ 0.105
XGBoost 1.549 $\pm$ 0.104 1.439 $\pm$ 0.104 1.443 $\pm$ 0.104
LightGBM 1.554 $\pm$ 0.093 1.480 $\pm$ 0.108 1.451 $\pm$ 0.108
CatBoost 1.583 $\pm$ 0.096 1.327 $\pm$ 0.101 1.330 $\pm$ 0.105
EBM 2.010 $\pm$ 0.117 1.955 $\pm$ 0.104 1.935 $\pm$ 0.113
FastAIMLP 2.327 $\pm$ 0.141 1.624 $\pm$ 0.100 1.646 $\pm$ 0.105
TorchMLP 1.493 $\pm$ 0.113 1.372 $\pm$ 0.105 1.374 $\pm$ 0.098
RealMLP 1.179 $\pm$ 0.085 1.146 $\pm$ 0.078 1.109 $\pm$ 0.081
TabM 1.249 $\pm$ 0.099 1.149 $\pm$ 0.092 1.140 $\pm$ 0.088
MNCA 1.539 $\pm$ 0.095 1.523 $\pm$ 0.137 1.452 $\pm$ 0.099
TabPFNv2 1.119 $\pm$ 0.088 1.112 $\pm$ 0.102 1.074 $\pm$ 0.094
TabDPT 1.203 $\pm$ 0.084 - -
TabICL - - -
Linear 5.344 $\pm$ 0.199 4.750 $\pm$ 0.140 4.743 $\pm$ 0.144
KNN 5.586 $\pm$ 0.184 5.458 $\pm$ 0.195 5.287 $\pm$ 0.163
AutoGluon - - 1.269 $\pm$ 0.090
anneal (logloss $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.046 $\pm$ 0.010 0.028 $\pm$ 0.025 0.023 $\pm$ 0.013
ExtraTrees 0.064 $\pm$ 0.012 0.028 $\pm$ 0.022 0.026 $\pm$ 0.022
XGBoost 0.039 $\pm$ 0.025 0.031 $\pm$ 0.025 0.031 $\pm$ 0.025
LightGBM 0.055 $\pm$ 0.026 0.033 $\pm$ 0.019 0.034 $\pm$ 0.019
CatBoost 0.040 $\pm$ 0.022 0.022 $\pm$ 0.021 0.021 $\pm$ 0.021
EBM 0.043 $\pm$ 0.025 0.036 $\pm$ 0.033 0.034 $\pm$ 0.032
FastAIMLP 0.085 $\pm$ 0.029 0.056 $\pm$ 0.022 0.054 $\pm$ 0.021
TorchMLP 0.040 $\pm$ 0.034 0.052 $\pm$ 0.044 0.040 $\pm$ 0.038
RealMLP 0.039 $\pm$ 0.031 0.029 $\pm$ 0.032 0.024 $\pm$ 0.026
TabM 0.036 $\pm$ 0.026 0.029 $\pm$ 0.027 0.028 $\pm$ 0.025
MNCA 0.043 $\pm$ 0.031 0.032 $\pm$ 0.038 0.033 $\pm$ 0.039
TabPFNv2 0.016 $\pm$ 0.014 0.023 $\pm$ 0.019 0.019 $\pm$ 0.014
TabDPT 0.058 $\pm$ 0.022 - -
TabICL 0.028 $\pm$ 0.014 - -
Linear 0.090 $\pm$ 0.028 0.047 $\pm$ 0.035 0.039 $\pm$ 0.028
KNN 1.064 $\pm$ 0.202 0.918 $\pm$ 0.209 0.531 $\pm$ 0.112
AutoGluon - - 0.037 $\pm$ 0.059
bank-marketing (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.726 $\pm$ 0.006 0.761 $\pm$ 0.005 0.761 $\pm$ 0.005
ExtraTrees 0.721 $\pm$ 0.004 0.758 $\pm$ 0.005 0.758 $\pm$ 0.005
XGBoost 0.763 $\pm$ 0.005 0.765 $\pm$ 0.006 0.765 $\pm$ 0.005
LightGBM 0.763 $\pm$ 0.005 0.765 $\pm$ 0.005 0.766 $\pm$ 0.005
CatBoost 0.766 $\pm$ 0.005 0.766 $\pm$ 0.005 0.765 $\pm$ 0.005
EBM 0.762 $\pm$ 0.005 0.763 $\pm$ 0.005 0.763 $\pm$ 0.005
FastAIMLP 0.759 $\pm$ 0.005 0.760 $\pm$ 0.006 0.761 $\pm$ 0.005
TorchMLP 0.757 $\pm$ 0.006 0.758 $\pm$ 0.006 0.759 $\pm$ 0.006
RealMLP 0.761 $\pm$ 0.005 0.763 $\pm$ 0.005 0.765 $\pm$ 0.005
TabM 0.764 $\pm$ 0.006 0.764 $\pm$ 0.005 0.765 $\pm$ 0.005
MNCA 0.762 $\pm$ 0.005 0.764 $\pm$ 0.006 0.764 $\pm$ 0.005
TabPFNv2 - - -
TabDPT 0.761 $\pm$ 0.005 - -
TabICL 0.764 $\pm$ 0.005 - -
Linear 0.748 $\pm$ 0.004 0.748 $\pm$ 0.004 0.748 $\pm$ 0.004
KNN 0.610 $\pm$ 0.005 0.650 $\pm$ 0.006 0.653 $\pm$ 0.007
AutoGluon - - 0.765 $\pm$ 0.006
blood-transfusion-service-center (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.682 $\pm$ 0.026 0.714 $\pm$ 0.029 0.713 $\pm$ 0.029
ExtraTrees 0.689 $\pm$ 0.025 0.728 $\pm$ 0.033 0.727 $\pm$ 0.031
XGBoost 0.708 $\pm$ 0.030 0.733 $\pm$ 0.031 0.731 $\pm$ 0.031
LightGBM 0.726 $\pm$ 0.033 0.743 $\pm$ 0.033 0.743 $\pm$ 0.031
CatBoost 0.738 $\pm$ 0.029 0.737 $\pm$ 0.031 0.736 $\pm$ 0.031
EBM 0.742 $\pm$ 0.032 0.743 $\pm$ 0.033 0.743 $\pm$ 0.033
FastAIMLP 0.743 $\pm$ 0.030 0.754 $\pm$ 0.030 0.756 $\pm$ 0.030
TorchMLP 0.749 $\pm$ 0.032 0.747 $\pm$ 0.030 0.748 $\pm$ 0.030
RealMLP 0.750 $\pm$ 0.030 0.737 $\pm$ 0.029 0.742 $\pm$ 0.027
TabM 0.741 $\pm$ 0.029 0.737 $\pm$ 0.031 0.741 $\pm$ 0.029
MNCA 0.756 $\pm$ 0.030 0.737 $\pm$ 0.032 0.715 $\pm$ 0.030
TabPFNv2 0.755 $\pm$ 0.029 0.748 $\pm$ 0.034 0.746 $\pm$ 0.030
TabDPT 0.751 $\pm$ 0.030 - -
TabICL 0.737 $\pm$ 0.031 - -
Linear 0.731 $\pm$ 0.032 0.747 $\pm$ 0.030 0.755 $\pm$ 0.026
KNN 0.661 $\pm$ 0.032 0.658 $\pm$ 0.036 0.680 $\pm$ 0.033
AutoGluon - - 0.748 $\pm$ 0.032
churn (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.915 $\pm$ 0.009 0.913 $\pm$ 0.012 0.913 $\pm$ 0.011
ExtraTrees 0.917 $\pm$ 0.011 0.919 $\pm$ 0.011 0.921 $\pm$ 0.011
XGBoost 0.923 $\pm$ 0.009 0.921 $\pm$ 0.011 0.920 $\pm$ 0.011
LightGBM 0.916 $\pm$ 0.011 0.920 $\pm$ 0.011 0.920 $\pm$ 0.011
CatBoost 0.924 $\pm$ 0.011 0.920 $\pm$ 0.012 0.922 $\pm$ 0.011
EBM 0.922 $\pm$ 0.014 0.924 $\pm$ 0.011 0.924 $\pm$ 0.011
FastAIMLP 0.918 $\pm$ 0.011 0.921 $\pm$ 0.009 0.921 $\pm$ 0.010
TorchMLP 0.888 $\pm$ 0.009 0.918 $\pm$ 0.010 0.918 $\pm$ 0.011
RealMLP 0.920 $\pm$ 0.010 0.924 $\pm$ 0.012 0.927 $\pm$ 0.011
TabM 0.925 $\pm$ 0.011 0.923 $\pm$ 0.010 0.923 $\pm$ 0.010
MNCA 0.903 $\pm$ 0.013 0.932 $\pm$ 0.013 0.931 $\pm$ 0.013
TabPFNv2 0.928 $\pm$ 0.011 0.925 $\pm$ 0.008 0.924 $\pm$ 0.010
TabDPT 0.923 $\pm$ 0.009 - -
TabICL 0.924 $\pm$ 0.011 - -
Linear 0.777 $\pm$ 0.018 0.816 $\pm$ 0.013 0.816 $\pm$ 0.013
KNN 0.681 $\pm$ 0.021 0.740 $\pm$ 0.018 0.739 $\pm$ 0.019
AutoGluon - - 0.922 $\pm$ 0.011
coil2000_insurance_policies (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.697 $\pm$ 0.014 0.741 $\pm$ 0.017 0.742 $\pm$ 0.016
ExtraTrees 0.696 $\pm$ 0.016 0.744 $\pm$ 0.016 0.748 $\pm$ 0.017
XGBoost 0.757 $\pm$ 0.015 0.757 $\pm$ 0.016 0.758 $\pm$ 0.014
LightGBM 0.752 $\pm$ 0.014 0.759 $\pm$ 0.015 0.761 $\pm$ 0.015
CatBoost 0.757 $\pm$ 0.014 0.758 $\pm$ 0.013 0.759 $\pm$ 0.012
EBM 0.754 $\pm$ 0.014 0.757 $\pm$ 0.013 0.761 $\pm$ 0.013
FastAIMLP 0.719 $\pm$ 0.010 0.749 $\pm$ 0.013 0.747 $\pm$ 0.013
TorchMLP 0.740 $\pm$ 0.011 0.747 $\pm$ 0.015 0.752 $\pm$ 0.014
RealMLP 0.742 $\pm$ 0.012 0.755 $\pm$ 0.011 0.763 $\pm$ 0.013
TabM 0.761 $\pm$ 0.012 0.763 $\pm$ 0.012 0.766 $\pm$ 0.012
MNCA 0.753 $\pm$ 0.013 0.757 $\pm$ 0.015 0.767 $\pm$ 0.012
TabPFNv2 0.753 $\pm$ 0.015 0.773 $\pm$ 0.015 0.773 $\pm$ 0.014
TabDPT 0.725 $\pm$ 0.010 - -
TabICL 0.756 $\pm$ 0.012 - -
Linear 0.737 $\pm$ 0.012 0.735 $\pm$ 0.011 0.736 $\pm$ 0.012
KNN 0.605 $\pm$ 0.010 0.676 $\pm$ 0.018 0.690 $\pm$ 0.009
AutoGluon - - 0.759 $\pm$ 0.016
concrete_compressive_strength (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 5.261 $\pm$ 0.336 5.189 $\pm$ 0.366 5.106 $\pm$ 0.352
ExtraTrees 5.139 $\pm$ 0.341 5.073 $\pm$ 0.333 5.048 $\pm$ 0.348
XGBoost 4.755 $\pm$ 0.387 4.236 $\pm$ 0.373 4.222 $\pm$ 0.384
LightGBM 4.484 $\pm$ 0.388 4.235 $\pm$ 0.395 4.212 $\pm$ 0.396
CatBoost 4.214 $\pm$ 0.413 4.231 $\pm$ 0.415 4.209 $\pm$ 0.411
EBM 4.442 $\pm$ 0.295 4.429 $\pm$ 0.331 4.371 $\pm$ 0.308
FastAIMLP 6.369 $\pm$ 0.379 5.187 $\pm$ 0.355 5.272 $\pm$ 0.360
TorchMLP 4.817 $\pm$ 0.354 4.715 $\pm$ 0.350 4.654 $\pm$ 0.334
RealMLP 4.688 $\pm$ 0.364 4.344 $\pm$ 0.289 4.133 $\pm$ 0.329
TabM 4.271 $\pm$ 0.385 4.269 $\pm$ 0.496 4.146 $\pm$ 0.409
MNCA 4.940 $\pm$ 0.309 4.749 $\pm$ 0.424 4.487 $\pm$ 0.388
TabPFNv2 4.259 $\pm$ 0.379 4.171 $\pm$ 0.439 4.118 $\pm$ 0.409
TabDPT 4.267 $\pm$ 0.422 - -
TabICL - - -
Linear 8.228 $\pm$ 0.359 8.159 $\pm$ 0.361 8.150 $\pm$ 0.354
KNN 9.556 $\pm$ 0.486 8.544 $\pm$ 0.533 8.419 $\pm$ 0.493
AutoGluon - - 4.165 $\pm$ 0.389
credit-g (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.783 $\pm$ 0.017 0.781 $\pm$ 0.019 0.782 $\pm$ 0.019
ExtraTrees 0.779 $\pm$ 0.019 0.781 $\pm$ 0.018 0.782 $\pm$ 0.018
XGBoost 0.783 $\pm$ 0.021 0.792 $\pm$ 0.021 0.793 $\pm$ 0.021
LightGBM 0.771 $\pm$ 0.019 0.792 $\pm$ 0.020 0.796 $\pm$ 0.020
CatBoost 0.789 $\pm$ 0.017 0.795 $\pm$ 0.020 0.795 $\pm$ 0.017
EBM 0.790 $\pm$ 0.021 0.782 $\pm$ 0.025 0.787 $\pm$ 0.023
FastAIMLP 0.783 $\pm$ 0.029 0.784 $\pm$ 0.025 0.793 $\pm$ 0.023
TorchMLP 0.772 $\pm$ 0.017 0.782 $\pm$ 0.020 0.788 $\pm$ 0.020
RealMLP 0.785 $\pm$ 0.022 0.784 $\pm$ 0.023 0.791 $\pm$ 0.020
TabM 0.793 $\pm$ 0.022 0.790 $\pm$ 0.020 0.795 $\pm$ 0.020
MNCA 0.783 $\pm$ 0.020 0.783 $\pm$ 0.020 0.775 $\pm$ 0.021
TabPFNv2 0.776 $\pm$ 0.019 0.773 $\pm$ 0.020 0.792 $\pm$ 0.020
TabDPT 0.780 $\pm$ 0.019 - -
TabICL 0.790 $\pm$ 0.017 - -
Linear 0.781 $\pm$ 0.021 0.780 $\pm$ 0.022 0.780 $\pm$ 0.021
KNN 0.558 $\pm$ 0.024 0.569 $\pm$ 0.027 0.579 $\pm$ 0.029
AutoGluon - - 0.794 $\pm$ 0.020
credit_card_clients_default (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.765 $\pm$ 0.005 0.780 $\pm$ 0.004 0.780 $\pm$ 0.004
ExtraTrees 0.766 $\pm$ 0.004 0.781 $\pm$ 0.004 0.782 $\pm$ 0.004
XGBoost 0.783 $\pm$ 0.004 0.785 $\pm$ 0.004 0.785 $\pm$ 0.004
LightGBM 0.784 $\pm$ 0.004 0.785 $\pm$ 0.004 0.785 $\pm$ 0.004
CatBoost 0.784 $\pm$ 0.004 0.785 $\pm$ 0.004 0.785 $\pm$ 0.004
EBM 0.783 $\pm$ 0.004 0.783 $\pm$ 0.004 0.784 $\pm$ 0.004
FastAIMLP 0.781 $\pm$ 0.005 0.783 $\pm$ 0.005 0.783 $\pm$ 0.005
TorchMLP 0.779 $\pm$ 0.003 0.783 $\pm$ 0.003 0.785 $\pm$ 0.003
RealMLP 0.785 $\pm$ 0.004 0.785 $\pm$ 0.005 0.786 $\pm$ 0.004
TabM 0.784 $\pm$ 0.004 0.788 $\pm$ 0.004 0.788 $\pm$ 0.004
MNCA 0.781 $\pm$ 0.004 0.787 $\pm$ 0.003 0.787 $\pm$ 0.004
TabPFNv2 - - -
TabDPT 0.780 $\pm$ 0.004 - -
TabICL 0.788 $\pm$ 0.004 - -
Linear 0.745 $\pm$ 0.004 0.745 $\pm$ 0.004 0.745 $\pm$ 0.004
KNN 0.605 $\pm$ 0.004 0.659 $\pm$ 0.005 0.666 $\pm$ 0.004
AutoGluon - - 0.787 $\pm$ 0.004
customer_satisfaction_in_airline (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.993 $\pm$ 0.000 0.993 $\pm$ 0.000 0.993 $\pm$ 0.000
ExtraTrees 0.992 $\pm$ 0.000 0.994 $\pm$ 0.000 0.994 $\pm$ 0.000
XGBoost 0.994 $\pm$ 0.000 0.994 $\pm$ 0.000 0.994 $\pm$ 0.000
LightGBM 0.994 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
CatBoost 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
EBM 0.985 $\pm$ 0.000 0.986 $\pm$ 0.001 0.986 $\pm$ 0.001
FastAIMLP 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
TorchMLP 0.993 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
RealMLP 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
TabM 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
MNCA 0.993 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
TabPFNv2 - - -
TabDPT 0.994 $\pm$ $0$
TabICL 0.995 $\pm$ $0$
Linear 0.964 $\pm$ 0.001 0.964 $\pm$ 0.001 0.965 $\pm$ 0.001
KNN 0.625 $\pm$ 0.002 0.676 $\pm$ 0.002 0.680 $\pm$ 0.002
AutoGluon - - 0.996 $\pm$ 0.000
diabetes (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.825 $\pm$ 0.023 0.830 $\pm$ 0.025 0.830 $\pm$ 0.024
ExtraTrees 0.826 $\pm$ 0.022 0.837 $\pm$ 0.021 0.837 $\pm$ 0.021
XGBoost 0.824 $\pm$ 0.024 0.830 $\pm$ 0.024 0.832 $\pm$ 0.024
LightGBM 0.829 $\pm$ 0.025 0.838 $\pm$ 0.026 0.837 $\pm$ 0.025
CatBoost 0.833 $\pm$ 0.025 0.834 $\pm$ 0.025 0.835 $\pm$ 0.024
EBM 0.840 $\pm$ 0.025 0.839 $\pm$ 0.023 0.840 $\pm$ 0.023
FastAIMLP 0.826 $\pm$ 0.024 0.832 $\pm$ 0.023 0.835 $\pm$ 0.023
TorchMLP 0.821 $\pm$ 0.024 0.825 $\pm$ 0.026 0.827 $\pm$ 0.025
RealMLP 0.833 $\pm$ 0.023 0.832 $\pm$ 0.022 0.836 $\pm$ 0.024
TabM 0.832 $\pm$ 0.024 0.830 $\pm$ 0.024 0.834 $\pm$ 0.024
MNCA 0.839 $\pm$ 0.024 0.834 $\pm$ 0.024 0.813 $\pm$ 0.022
TabPFNv2 0.844 $\pm$ 0.023 0.842 $\pm$ 0.024 0.839 $\pm$ 0.024
TabDPT 0.840 $\pm$ 0.023 - -
TabICL 0.837 $\pm$ 0.023 - -
Linear 0.832 $\pm$ 0.024 0.831 $\pm$ 0.023 0.831 $\pm$ 0.023
KNN 0.740 $\pm$ 0.030 0.809 $\pm$ 0.026 0.806 $\pm$ 0.027
AutoGluon - - 0.835 $\pm$ 0.023
diamonds (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 549.9 $\pm$ 8.3 549.9 $\pm$ 8.3 547.2 $\pm$ 9.3
ExtraTrees 536.6 $\pm$ 8.7 536.3 $\pm$ 9.1 534.6 $\pm$ 8.9
XGBoost 539.0 $\pm$ 10.0 530.1 $\pm$ 10.0 528.2 $\pm$ 10.9
LightGBM 532.1 $\pm$ 9.1 524.9 $\pm$ 9.7 519.0 $\pm$ 9.4
CatBoost 520.7 $\pm$ 12.0 520.7 $\pm$ 12.0 520.8 $\pm$ 11.8
EBM 618.2 $\pm$ 14.5 613.7 $\pm$ 11.9 612.4 $\pm$ 13.9
FastAIMLP 563.1 $\pm$ 15.0 559.4 $\pm$ 9.4 550.0 $\pm$ 8.9
TorchMLP 627.3 $\pm$ 25.8 550.5 $\pm$ 16.6 542.6 $\pm$ 15.2
RealMLP 529.6 $\pm$ 8.1 521.5 $\pm$ 7.6 513.7 $\pm$ 7.3
TabM 522.5 $\pm$ 8.9 521.5 $\pm$ 8.1 518.9 $\pm$ 8.8
MNCA 524.7 $\pm$ 8.7 521.1 $\pm$ 7.5 510.3 $\pm$ 8.1
TabPFNv2 - - -
TabDPT 535.4 $\pm$ 13.1 - -
TabICL - - -
Linear 1652.1 $\pm$ 177.2 1142.9 $\pm$ 28.3 1144.1 $\pm$ 29.7
KNN 1461.6 $\pm$ 19.2 1368.3 $\pm$ 19.4 1362.5 $\pm$ 19.5
AutoGluon - - 510.5 $\pm$ 9.5
hazelnut-spread-contaminant-detection (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.958 $\pm$ 0.006 0.959 $\pm$ 0.006 0.960 $\pm$ 0.006
ExtraTrees 0.955 $\pm$ 0.006 0.964 $\pm$ 0.005 0.964 $\pm$ 0.005
XGBoost 0.973 $\pm$ 0.005 0.975 $\pm$ 0.004 0.975 $\pm$ 0.004
LightGBM 0.973 $\pm$ 0.005 0.978 $\pm$ 0.004 0.978 $\pm$ 0.004
CatBoost 0.974 $\pm$ 0.004 0.975 $\pm$ 0.004 0.974 $\pm$ 0.004
EBM 0.971 $\pm$ 0.005 0.975 $\pm$ 0.004 0.976 $\pm$ 0.004
FastAIMLP 0.983 $\pm$ 0.003 0.986 $\pm$ 0.003 0.986 $\pm$ 0.003
TorchMLP 0.983 $\pm$ 0.003 0.987 $\pm$ 0.003 0.987 $\pm$ 0.003
RealMLP 0.984 $\pm$ 0.003 0.986 $\pm$ 0.003 0.986 $\pm$ 0.003
TabM 0.967 $\pm$ 0.005 0.984 $\pm$ 0.003 0.985 $\pm$ 0.003
MNCA 0.986 $\pm$ 0.003 0.988 $\pm$ 0.003 0.988 $\pm$ 0.003
TabPFNv2 0.988 $\pm$ 0.003 0.989 $\pm$ 0.003 0.989 $\pm$ 0.003
TabDPT 0.992 $\pm$ 0.002 - -
TabICL 0.992 $\pm$ 0.002 - -
Linear 0.948 $\pm$ 0.006 0.951 $\pm$ 0.006 0.952 $\pm$ 0.006
KNN 0.908 $\pm$ 0.009 0.922 $\pm$ 0.010 0.931 $\pm$ 0.008
AutoGluon - - 0.987 $\pm$ 0.003
healthcare_insurance_expenses (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 4889.0 $\pm$ 289.0 4641.0 $\pm$ 304.0 4630.0 $\pm$ 303.0
ExtraTrees 4845.0 $\pm$ 275.0 4607.0 $\pm$ 324.0 4610.0 $\pm$ 323.0
XGBoost 4672.0 $\pm$ 306.0 4523.0 $\pm$ 320.0 4520.0 $\pm$ 320.0
LightGBM 4610.0 $\pm$ 314.0 4525.0 $\pm$ 329.0 4512.0 $\pm$ 325.0
CatBoost 4535.0 $\pm$ 329.0 4518.0 $\pm$ 322.0 4519.0 $\pm$ 321.0
EBM 4549.0 $\pm$ 319.0 4500.0 $\pm$ 333.0 4499.0 $\pm$ 326.0
FastAIMLP 4720.0 $\pm$ 313.0 4634.0 $\pm$ 318.0 4624.0 $\pm$ 311.0
TorchMLP 4661.0 $\pm$ 343.0 4526.0 $\pm$ 329.0 4534.0 $\pm$ 333.0
RealMLP 4579.0 $\pm$ 313.0 4571.0 $\pm$ 324.0 4535.0 $\pm$ 323.0
TabM 4514.0 $\pm$ 323.0 4530.0 $\pm$ 323.0 4511.0 $\pm$ 326.0
MNCA 4605.0 $\pm$ 332.0 4614.0 $\pm$ 330.0 4589.0 $\pm$ 331.0
TabPFNv2 4695.0 $\pm$ 303.0 4650.0 $\pm$ 337.0 4568.0 $\pm$ 319.0
TabDPT 4508.0 $\pm$ 295.0 - -
TabICL - - -
Linear 6083.0 $\pm$ 276.0 6085.0 $\pm$ 276.0 6085.0 $\pm$ 276.0
KNN 12371.0 $\pm$ 504.0 11514.0 $\pm$ 472.0 11540.0 $\pm$ 478.0
AutoGluon - - 4490.0 $\pm$ 332.0
heloc (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.791 $\pm$ 0.005 0.792 $\pm$ 0.006 0.793 $\pm$ 0.005
ExtraTrees 0.790 $\pm$ 0.005 0.793 $\pm$ 0.005 0.793 $\pm$ 0.005
XGBoost 0.794 $\pm$ 0.005 0.797 $\pm$ 0.005 0.797 $\pm$ 0.005
LightGBM 0.794 $\pm$ 0.005 0.799 $\pm$ 0.005 0.799 $\pm$ 0.005
CatBoost 0.798 $\pm$ 0.004 0.798 $\pm$ 0.005 0.798 $\pm$ 0.004
EBM 0.799 $\pm$ 0.005 0.799 $\pm$ 0.005 0.799 $\pm$ 0.005
FastAIMLP 0.791 $\pm$ 0.004 0.794 $\pm$ 0.005 0.795 $\pm$ 0.005
TorchMLP 0.791 $\pm$ 0.004 0.795 $\pm$ 0.004 0.796 $\pm$ 0.004
RealMLP 0.798 $\pm$ 0.004 0.798 $\pm$ 0.004 0.800 $\pm$ 0.004
TabM 0.797 $\pm$ 0.004 0.799 $\pm$ 0.004 0.799 $\pm$ 0.004
MNCA 0.799 $\pm$ 0.004 0.801 $\pm$ 0.005 0.799 $\pm$ 0.005
TabPFNv2 0.801 $\pm$ 0.003 0.801 $\pm$ 0.004 0.801 $\pm$ 0.003
TabDPT 0.794 $\pm$ 0.004 - -
TabICL 0.800 $\pm$ 0.004 - -
Linear 0.786 $\pm$ 0.005 0.786 $\pm$ 0.005 0.786 $\pm$ 0.005
KNN 0.691 $\pm$ 0.007 0.747 $\pm$ 0.003 0.747 $\pm$ 0.003
AutoGluon - - 0.798 $\pm$ 0.005
hiva_agnostic (logloss $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.263 $\pm$ 0.025 0.174 $\pm$ 0.001 0.174 $\pm$ 0.001
ExtraTrees 0.268 $\pm$ 0.027 0.174 $\pm$ 0.000 0.174 $\pm$ 0.000
XGBoost 0.182 $\pm$ 0.002 0.179 $\pm$ 0.002 0.179 $\pm$ 0.002
LightGBM 0.175 $\pm$ 0.001 0.175 $\pm$ 0.001 0.175 $\pm$ 0.001
CatBoost 0.176 $\pm$ 0.001 0.177 $\pm$ 0.002 0.177 $\pm$ 0.002
EBM 0.174 $\pm$ 0.001 0.176 $\pm$ 0.001 0.175 $\pm$ 0.001
FastAIMLP 0.213 $\pm$ 0.010 0.183 $\pm$ 0.008 0.183 $\pm$ 0.004
TorchMLP 0.183 $\pm$ 0.005 0.176 $\pm$ 0.001 0.178 $\pm$ 0.003
RealMLP 0.196 $\pm$ 0.007 0.176 $\pm$ 0.002 0.179 $\pm$ 0.002
TabM 0.177 $\pm$ 0.001 0.175 $\pm$ 0.001 0.175 $\pm$ 0.001
MNCA 0.224 $\pm$ 0.012 0.176 $\pm$ 0.002 0.179 $\pm$ 0.002
TabPFNv2 - - -
TabDPT 0.181 $\pm$ 0.004 - -
TabICL - - -
Linear 0.448 $\pm$ 0.024 0.449 $\pm$ 0.024 0.450 $\pm$ 0.024
KNN 0.263 $\pm$ 0.025 0.264 $\pm$ 0.026 0.264 $\pm$ 0.026
AutoGluon - - 0.193 $\pm$ 0.027
houses (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.231 $\pm$ 0.002 0.231 $\pm$ 0.002 0.230 $\pm$ 0.002
ExtraTrees 0.243 $\pm$ 0.002 0.238 $\pm$ 0.002 0.238 $\pm$ 0.002
XGBoost 0.215 $\pm$ 0.003 0.215 $\pm$ 0.002 0.215 $\pm$ 0.002
LightGBM 0.217 $\pm$ 0.002 0.212 $\pm$ 0.002 0.211 $\pm$ 0.002
CatBoost 0.211 $\pm$ 0.002 0.211 $\pm$ 0.002 0.211 $\pm$ 0.002
EBM 0.231 $\pm$ 0.003 0.229 $\pm$ 0.003 0.228 $\pm$ 0.003
FastAIMLP 0.244 $\pm$ 0.001 0.236 $\pm$ 0.002 0.235 $\pm$ 0.001
TorchMLP 0.233 $\pm$ 0.003 0.228 $\pm$ 0.003 0.226 $\pm$ 0.002
RealMLP 0.223 $\pm$ 0.002 0.211 $\pm$ 0.003 0.203 $\pm$ 0.003
TabM 0.212 $\pm$ 0.002 0.208 $\pm$ 0.002 0.205 $\pm$ 0.002
MNCA 0.204 $\pm$ 0.003 0.203 $\pm$ 0.003 0.199 $\pm$ 0.002
TabPFNv2 - - -
TabDPT 0.209 $\pm$ 0.003 - -
TabICL - - -
Linear 0.325 $\pm$ 0.003 0.325 $\pm$ 0.003 0.324 $\pm$ 0.003
KNN 0.516 $\pm$ 0.004 0.480 $\pm$ 0.004 0.478 $\pm$ 0.004
AutoGluon - - 0.204 $\pm$ 0.002
in_vehicle_coupon_recommendation (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.812 $\pm$ 0.007 0.817 $\pm$ 0.008 0.822 $\pm$ 0.008
ExtraTrees 0.798 $\pm$ 0.007 0.804 $\pm$ 0.008 0.811 $\pm$ 0.008
XGBoost 0.832 $\pm$ 0.004 0.842 $\pm$ 0.005 0.843 $\pm$ 0.005
LightGBM 0.836 $\pm$ 0.005 0.844 $\pm$ 0.005 0.845 $\pm$ 0.005
CatBoost 0.840 $\pm$ 0.006 0.843 $\pm$ 0.005 0.844 $\pm$ 0.006
EBM 0.802 $\pm$ 0.006 0.807 $\pm$ 0.006 0.807 $\pm$ 0.006
FastAIMLP 0.810 $\pm$ 0.007 0.823 $\pm$ 0.007 0.826 $\pm$ 0.006
TorchMLP 0.825 $\pm$ 0.004 0.833 $\pm$ 0.008 0.841 $\pm$ 0.006
RealMLP 0.837 $\pm$ 0.006 0.839 $\pm$ 0.006 0.849 $\pm$ 0.006
TabM 0.848 $\pm$ 0.005 0.851 $\pm$ 0.006 0.852 $\pm$ 0.006
MNCA 0.812 $\pm$ 0.006 0.842 $\pm$ 0.006 0.849 $\pm$ 0.006
TabPFNv2 0.789 $\pm$ 0.008 0.806 $\pm$ 0.008 0.837 $\pm$ 0.007
TabDPT 0.798 $\pm$ 0.005 - -
TabICL 0.846 $\pm$ 0.006 - -
Linear 0.735 $\pm$ 0.007 0.735 $\pm$ 0.007 0.735 $\pm$ 0.007
KNN 0.500 $\pm$ 0.000 0.502 $\pm$ 0.005 0.502 $\pm$ 0.005
AutoGluon - - 0.847 $\pm$ 0.006
jm1 (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.752 $\pm$ 0.008 0.752 $\pm$ 0.008 0.761 $\pm$ 0.007
ExtraTrees 0.756 $\pm$ 0.007 0.758 $\pm$ 0.008 0.765 $\pm$ 0.006
XGBoost 0.748 $\pm$ 0.007 0.749 $\pm$ 0.007 0.752 $\pm$ 0.006
LightGBM 0.748 $\pm$ 0.006 0.751 $\pm$ 0.006 0.753 $\pm$ 0.006
CatBoost 0.744 $\pm$ 0.005 0.751 $\pm$ 0.005 0.749 $\pm$ 0.005
EBM 0.735 $\pm$ 0.006 0.734 $\pm$ 0.007 0.736 $\pm$ 0.007
FastAIMLP 0.728 $\pm$ 0.007 0.728 $\pm$ 0.007 0.733 $\pm$ 0.006
TorchMLP 0.728 $\pm$ 0.005 0.734 $\pm$ 0.005 0.736 $\pm$ 0.005
RealMLP 0.731 $\pm$ 0.007 0.735 $\pm$ 0.007 0.749 $\pm$ 0.007
TabM 0.733 $\pm$ 0.007 0.738 $\pm$ 0.004 0.746 $\pm$ 0.006
MNCA 0.762 $\pm$ 0.003 0.761 $\pm$ 0.006 0.769 $\pm$ 0.006
TabPFNv2 0.732 $\pm$ 0.008 0.755 $\pm$ 0.007 0.773 $\pm$ 0.006
TabDPT 0.771 $\pm$ 0.005 - -
TabICL 0.776 $\pm$ 0.005 - -
Linear 0.724 $\pm$ 0.006 0.724 $\pm$ 0.006 0.724 $\pm$ 0.007
KNN 0.648 $\pm$ 0.008 0.730 $\pm$ 0.009 0.730 $\pm$ 0.009
AutoGluon - - 0.760 $\pm$ 0.007
kddcup09_appetency (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.772 $\pm$ 0.016 0.822 $\pm$ 0.011 0.821 $\pm$ 0.010
ExtraTrees 0.771 $\pm$ 0.012 0.819 $\pm$ 0.012 0.821 $\pm$ 0.009
XGBoost 0.830 $\pm$ 0.012 0.833 $\pm$ 0.009 0.837 $\pm$ 0.010
LightGBM 0.798 $\pm$ 0.009 0.821 $\pm$ 0.009 0.829 $\pm$ 0.010
CatBoost 0.846 $\pm$ 0.008 0.845 $\pm$ 0.008 0.845 $\pm$ 0.008
EBM 0.826 $\pm$ 0.009 0.831 $\pm$ 0.010 0.833 $\pm$ 0.010
FastAIMLP 0.749 $\pm$ 0.023 0.795 $\pm$ 0.013 0.804 $\pm$ 0.014
TorchMLP 0.819 $\pm$ 0.012 0.826 $\pm$ 0.012 0.831 $\pm$ 0.013
RealMLP 0.820 $\pm$ 0.011 0.822 $\pm$ 0.011 0.832 $\pm$ 0.011
TabM 0.806 $\pm$ 0.011 0.821 $\pm$ 0.010 0.821 $\pm$ 0.010
MNCA 0.797 $\pm$ 0.010 0.814 $\pm$ 0.012 0.803 $\pm$ 0.015
TabPFNv2 - - -
TabDPT 0.742 $\pm$ 0.009 - -
TabICL 0.811 $\pm$ 0.014 - -
Linear 0.797 $\pm$ 0.013 0.797 $\pm$ 0.013 0.796 $\pm$ 0.014
KNN 0.511 $\pm$ 0.007 0.553 $\pm$ 0.010 0.551 $\pm$ 0.011
AutoGluon - - 0.846 $\pm$ 0.009
maternal_health_risk (logloss $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.479 $\pm$ 0.071 0.470 $\pm$ 0.053 0.448 $\pm$ 0.054
ExtraTrees 0.478 $\pm$ 0.069 0.453 $\pm$ 0.055 0.443 $\pm$ 0.055
XGBoost 0.470 $\pm$ 0.051 0.459 $\pm$ 0.051 0.462 $\pm$ 0.050
LightGBM 0.488 $\pm$ 0.052 0.462 $\pm$ 0.049 0.461 $\pm$ 0.045
CatBoost 0.478 $\pm$ 0.055 0.463 $\pm$ 0.053 0.459 $\pm$ 0.049
EBM 0.569 $\pm$ 0.038 0.562 $\pm$ 0.042 0.557 $\pm$ 0.040
FastAIMLP 0.650 $\pm$ 0.044 0.617 $\pm$ 0.040 0.611 $\pm$ 0.039
TorchMLP 0.606 $\pm$ 0.054 0.566 $\pm$ 0.053 0.554 $\pm$ 0.046
RealMLP 0.558 $\pm$ 0.056 0.463 $\pm$ 0.058 0.436 $\pm$ 0.049
TabM 0.513 $\pm$ 0.045 0.484 $\pm$ 0.057 0.469 $\pm$ 0.051
MNCA 0.453 $\pm$ 0.042 0.442 $\pm$ 0.047 0.428 $\pm$ 0.050
TabPFNv2 0.451 $\pm$ 0.047 0.439 $\pm$ 0.057 0.437 $\pm$ 0.057
TabDPT 0.405 $\pm$ 0.062 - -
TabICL 0.410 $\pm$ 0.058 - -
Linear 0.796 $\pm$ 0.037 0.795 $\pm$ 0.042 0.784 $\pm$ 0.039
KNN 1.372 $\pm$ 0.265 0.869 $\pm$ 0.149 0.490 $\pm$ 0.062
AutoGluon - - 0.462 $\pm$ 0.061
miami_housing (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 9676 $\pm$ 592 9332 $\pm$ 503 9302 $\pm$ 500
ExtraTrees 9482 $\pm$ 400 9195 $\pm$ 395 9166 $\pm$ 409
XGBoost 8650 $\pm$ 447 8062 $\pm$ 361 8042 $\pm$ 357
LightGBM 8563 $\pm$ 463 8124 $\pm$ 410 7961 $\pm$ 354
CatBoost 7985 $\pm$ 315 7836 $\pm$ 342 7847 $\pm$ 329
EBM 10420 $\pm$ 365 9882 $\pm$ 466 9823 $\pm$ 438
FastAIMLP 9034 $\pm$ 512 8855 $\pm$ 535 8664 $\pm$ 483
TorchMLP 9265 $\pm$ 455 8631 $\pm$ 511 8528 $\pm$ 466
RealMLP 8605 $\pm$ 402 8337 $\pm$ 457 8018 $\pm$ 399
TabM 8307 $\pm$ 486 8115 $\pm$ 449 8023 $\pm$ 428
MNCA 8813 $\pm$ 434 8307 $\pm$ 380 8015 $\pm$ 409
TabPFNv2 8579 $\pm$ 447 7829 $\pm$ 457 7711 $\pm$ 442
TabDPT 8213 $\pm$ 497 - -
TabICL - - -
Linear 19126 $\pm$ 458 17554 $\pm$ 363 17554 $\pm$ 363
KNN 14211 $\pm$ 438 13300 $\pm$ 442 13148 $\pm$ 399
AutoGluon - - 7873 $\pm$ 429
online_shoppers_intention (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.926 $\pm$ 0.004 0.931 $\pm$ 0.004 0.932 $\pm$ 0.003
ExtraTrees 0.918 $\pm$ 0.005 0.931 $\pm$ 0.004 0.931 $\pm$ 0.004
XGBoost 0.935 $\pm$ 0.004 0.934 $\pm$ 0.003 0.936 $\pm$ 0.003
LightGBM 0.934 $\pm$ 0.003 0.935 $\pm$ 0.004 0.935 $\pm$ 0.003
CatBoost 0.934 $\pm$ 0.003 0.933 $\pm$ 0.004 0.934 $\pm$ 0.004
EBM 0.931 $\pm$ 0.003 0.931 $\pm$ 0.003 0.931 $\pm$ 0.003
FastAIMLP 0.926 $\pm$ 0.004 0.931 $\pm$ 0.005 0.932 $\pm$ 0.004
TorchMLP 0.930 $\pm$ 0.004 0.935 $\pm$ 0.004 0.936 $\pm$ 0.003
RealMLP 0.929 $\pm$ 0.004 0.933 $\pm$ 0.003 0.934 $\pm$ 0.004
TabM 0.935 $\pm$ 0.003 0.936 $\pm$ 0.003 0.936 $\pm$ 0.004
MNCA 0.934 $\pm$ 0.003 0.935 $\pm$ 0.003 0.936 $\pm$ 0.003
TabPFNv2 0.934 $\pm$ 0.004 0.937 $\pm$ 0.003 0.937 $\pm$ 0.003
TabDPT 0.926 $\pm$ 0.005 - -
TabICL 0.937 $\pm$ 0.003 - -
Linear 0.913 $\pm$ 0.007 0.913 $\pm$ 0.007 0.917 $\pm$ 0.007
KNN 0.759 $\pm$ 0.009 0.828 $\pm$ 0.006 0.832 $\pm$ 0.003
AutoGluon - - 0.936 $\pm$ 0.003
physiochemical_protein (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 3.565 $\pm$ 0.021 3.463 $\pm$ 0.021 3.471 $\pm$ 0.021
ExtraTrees 3.548 $\pm$ 0.022 3.466 $\pm$ 0.022 3.474 $\pm$ 0.021
XGBoost 3.513 $\pm$ 0.024 3.390 $\pm$ 0.024 3.390 $\pm$ 0.024
LightGBM 3.477 $\pm$ 0.026 3.381 $\pm$ 0.027 3.384 $\pm$ 0.027
CatBoost 3.522 $\pm$ 0.026 3.395 $\pm$ 0.029 3.383 $\pm$ 0.027
EBM 4.241 $\pm$ 0.020 4.234 $\pm$ 0.019 4.226 $\pm$ 0.020
FastAIMLP 4.014 $\pm$ 0.027 3.675 $\pm$ 0.031 3.683 $\pm$ 0.034
TorchMLP 3.388 $\pm$ 0.021 3.289 $\pm$ 0.028 3.228 $\pm$ 0.018
RealMLP 3.466 $\pm$ 0.042 3.284 $\pm$ 0.029 3.125 $\pm$ 0.028
TabM 3.426 $\pm$ 0.027 3.273 $\pm$ 0.031 3.212 $\pm$ 0.032
MNCA 3.170 $\pm$ 0.040 3.056 $\pm$ 0.034 2.985 $\pm$ 0.038
TabPFNv2 - - -
TabDPT 2.912 $\pm$ 0.033 - -
TabICL - - -
Linear 5.194 $\pm$ 0.023 5.188 $\pm$ 0.022 5.142 $\pm$ 0.023
KNN 5.955 $\pm$ 0.028 5.482 $\pm$ 0.027 5.453 $\pm$ 0.026
AutoGluon - - 3.107 $\pm$ 0.026
polish_companies_bankruptcy (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.927 $\pm$ 0.007 0.935 $\pm$ 0.008 0.938 $\pm$ 0.007
ExtraTrees 0.874 $\pm$ 0.013 0.891 $\pm$ 0.013 0.893 $\pm$ 0.011
XGBoost 0.958 $\pm$ 0.007 0.957 $\pm$ 0.007 0.957 $\pm$ 0.008
LightGBM 0.954 $\pm$ 0.007 0.955 $\pm$ 0.008 0.957 $\pm$ 0.007
CatBoost 0.961 $\pm$ 0.008 0.961 $\pm$ 0.008 0.960 $\pm$ 0.008
EBM 0.962 $\pm$ 0.009 0.962 $\pm$ 0.010 0.964 $\pm$ 0.009
FastAIMLP 0.841 $\pm$ 0.026 0.852 $\pm$ 0.034 0.862 $\pm$ 0.022
TorchMLP 0.903 $\pm$ 0.006 0.955 $\pm$ 0.006 0.957 $\pm$ 0.005
RealMLP 0.962 $\pm$ 0.004 0.957 $\pm$ 0.006 0.963 $\pm$ 0.006
TabM 0.951 $\pm$ 0.009 0.969 $\pm$ 0.005 0.970 $\pm$ 0.004
MNCA 0.962 $\pm$ 0.008 0.965 $\pm$ 0.009 0.968 $\pm$ 0.006
TabPFNv2 0.959 $\pm$ 0.006 0.979 $\pm$ 0.003 0.981 $\pm$ 0.002
TabDPT 0.958 $\pm$ 0.009 - -
TabICL 0.974 $\pm$ 0.002 - -
Linear 0.867 $\pm$ 0.016 0.887 $\pm$ 0.013 0.891 $\pm$ 0.013
KNN 0.698 $\pm$ 0.015 0.784 $\pm$ 0.015 0.786 $\pm$ 0.017
AutoGluon - - 0.969 $\pm$ 0.005
qsar-biodeg (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.930 $\pm$ 0.013 0.929 $\pm$ 0.013 0.929 $\pm$ 0.013
ExtraTrees 0.932 $\pm$ 0.013 0.932 $\pm$ 0.013 0.934 $\pm$ 0.013
XGBoost 0.926 $\pm$ 0.013 0.931 $\pm$ 0.012 0.931 $\pm$ 0.012
LightGBM 0.927 $\pm$ 0.012 0.933 $\pm$ 0.012 0.933 $\pm$ 0.012
CatBoost 0.930 $\pm$ 0.012 0.931 $\pm$ 0.012 0.932 $\pm$ 0.011
EBM 0.931 $\pm$ 0.011 0.931 $\pm$ 0.011 0.933 $\pm$ 0.011
FastAIMLP 0.932 $\pm$ 0.013 0.932 $\pm$ 0.014 0.934 $\pm$ 0.013
TorchMLP 0.924 $\pm$ 0.014 0.923 $\pm$ 0.014 0.927 $\pm$ 0.014
RealMLP 0.927 $\pm$ 0.013 0.926 $\pm$ 0.015 0.934 $\pm$ 0.012
TabM 0.931 $\pm$ 0.011 0.934 $\pm$ 0.013 0.936 $\pm$ 0.012
MNCA 0.928 $\pm$ 0.012 0.928 $\pm$ 0.012 0.931 $\pm$ 0.012
TabPFNv2 0.936 $\pm$ 0.011 0.932 $\pm$ 0.013 0.936 $\pm$ 0.012
TabDPT 0.934 $\pm$ 0.012 - -
TabICL 0.938 $\pm$ 0.012 - -
Linear 0.910 $\pm$ 0.016 0.917 $\pm$ 0.014 0.918 $\pm$ 0.014
KNN 0.862 $\pm$ 0.018 0.894 $\pm$ 0.022 0.898 $\pm$ 0.019
AutoGluon - - 0.934 $\pm$ 0.013
seismic-bumps (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.747 $\pm$ 0.021 0.767 $\pm$ 0.022 0.765 $\pm$ 0.026
ExtraTrees 0.734 $\pm$ 0.024 0.767 $\pm$ 0.029 0.771 $\pm$ 0.025
XGBoost 0.759 $\pm$ 0.022 0.768 $\pm$ 0.024 0.771 $\pm$ 0.025
LightGBM 0.752 $\pm$ 0.027 0.770 $\pm$ 0.027 0.771 $\pm$ 0.026
CatBoost 0.776 $\pm$ 0.027 0.772 $\pm$ 0.026 0.767 $\pm$ 0.028
EBM 0.770 $\pm$ 0.026 0.763 $\pm$ 0.026 0.767 $\pm$ 0.025
FastAIMLP 0.728 $\pm$ 0.034 0.759 $\pm$ 0.033 0.761 $\pm$ 0.028
TorchMLP 0.763 $\pm$ 0.026 0.758 $\pm$ 0.026 0.762 $\pm$ 0.025
RealMLP 0.760 $\pm$ 0.030 0.761 $\pm$ 0.027 0.766 $\pm$ 0.027
TabM 0.769 $\pm$ 0.024 0.768 $\pm$ 0.026 0.770 $\pm$ 0.026
MNCA 0.771 $\pm$ 0.023 0.763 $\pm$ 0.027 0.734 $\pm$ 0.026
TabPFNv2 0.772 $\pm$ 0.025 0.766 $\pm$ 0.023 0.769 $\pm$ 0.024
TabDPT 0.774 $\pm$ 0.022 - -
TabICL 0.783 $\pm$ 0.024 - -
Linear 0.759 $\pm$ 0.024 0.757 $\pm$ 0.023 0.761 $\pm$ 0.025
KNN 0.594 $\pm$ 0.019 0.702 $\pm$ 0.026 0.701 $\pm$ 0.031
AutoGluon - - 0.758 $\pm$ 0.032
splice (logloss $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.317 $\pm$ 0.005 0.180 $\pm$ 0.015 0.178 $\pm$ 0.014
ExtraTrees 0.393 $\pm$ 0.007 0.175 $\pm$ 0.015 0.173 $\pm$ 0.013
XGBoost 0.119 $\pm$ 0.020 0.109 $\pm$ 0.018 0.109 $\pm$ 0.018
LightGBM 0.111 $\pm$ 0.016 0.103 $\pm$ 0.016 0.102 $\pm$ 0.015
CatBoost 0.110 $\pm$ 0.017 0.115 $\pm$ 0.020 0.111 $\pm$ 0.018
EBM 0.118 $\pm$ 0.013 0.119 $\pm$ 0.012 0.117 $\pm$ 0.012
FastAIMLP 0.118 $\pm$ 0.013 0.105 $\pm$ 0.013 0.103 $\pm$ 0.014
TorchMLP 0.157 $\pm$ 0.022 0.127 $\pm$ 0.017 0.116 $\pm$ 0.014
RealMLP 0.126 $\pm$ 0.014 0.110 $\pm$ 0.014 0.106 $\pm$ 0.013
TabM 0.110 $\pm$ 0.016 0.112 $\pm$ 0.019 0.110 $\pm$ 0.017
MNCA 0.144 $\pm$ 0.013 0.120 $\pm$ 0.011 0.122 $\pm$ 0.014
TabPFNv2 0.107 $\pm$ 0.015 0.113 $\pm$ 0.019 0.099 $\pm$ 0.015
TabDPT 0.267 $\pm$ 0.010 - -
TabICL 0.148 $\pm$ 0.020 - -
Linear 0.167 $\pm$ 0.019 0.167 $\pm$ 0.019 0.167 $\pm$ 0.019
KNN 0.317 $\pm$ 0.005 0.317 $\pm$ 0.005 0.317 $\pm$ 0.005
AutoGluon - - 0.100 $\pm$ 0.017
students_dropout_and_academic_success (logloss $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.584 $\pm$ 0.011 0.576 $\pm$ 0.014 0.574 $\pm$ 0.013
ExtraTrees 0.591 $\pm$ 0.012 0.570 $\pm$ 0.011 0.566 $\pm$ 0.013
XGBoost 0.554 $\pm$ 0.016 0.545 $\pm$ 0.015 0.546 $\pm$ 0.014
LightGBM 0.555 $\pm$ 0.015 0.543 $\pm$ 0.016 0.542 $\pm$ 0.015
CatBoost 0.552 $\pm$ 0.016 0.543 $\pm$ 0.017 0.541 $\pm$ 0.016
EBM 0.565 $\pm$ 0.014 0.561 $\pm$ 0.013 0.560 $\pm$ 0.014
FastAIMLP 0.565 $\pm$ 0.021 0.549 $\pm$ 0.015 0.540 $\pm$ 0.015
TorchMLP 0.581 $\pm$ 0.018 0.559 $\pm$ 0.016 0.552 $\pm$ 0.014
RealMLP 0.556 $\pm$ 0.015 0.553 $\pm$ 0.013 0.542 $\pm$ 0.014
TabM 0.543 $\pm$ 0.013 0.541 $\pm$ 0.014 0.538 $\pm$ 0.014
MNCA 0.555 $\pm$ 0.015 0.554 $\pm$ 0.012 0.546 $\pm$ 0.013
TabPFNv2 0.534 $\pm$ 0.012 0.529 $\pm$ 0.015 0.527 $\pm$ 0.015
TabDPT 0.561 $\pm$ 0.018 - -
TabICL 0.550 $\pm$ 0.014 - -
Linear 0.571 $\pm$ 0.017 0.571 $\pm$ 0.017 0.571 $\pm$ 0.016
KNN 1.957 $\pm$ 0.115 0.721 $\pm$ 0.005 0.715 $\pm$ 0.006
AutoGluon - - 0.536 $\pm$ 0.015
superconductivity (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 9.63 $\pm$ 0.19 9.62 $\pm$ 0.19 9.53 $\pm$ 0.19
ExtraTrees 9.43 $\pm$ 0.18 9.43 $\pm$ 0.18 9.39 $\pm$ 0.18
XGBoost 9.45 $\pm$ 0.18 9.35 $\pm$ 0.21 9.34 $\pm$ 0.20
LightGBM 9.41 $\pm$ 0.21 9.28 $\pm$ 0.22 9.26 $\pm$ 0.20
CatBoost 9.36 $\pm$ 0.19 9.34 $\pm$ 0.21 9.34 $\pm$ 0.20
EBM 10.53 $\pm$ 0.23 10.43 $\pm$ 0.22 10.21 $\pm$ 0.18
FastAIMLP 11.51 $\pm$ 0.10 10.59 $\pm$ 0.10 10.55 $\pm$ 0.12
TorchMLP 9.95 $\pm$ 0.21 9.66 $\pm$ 0.16 9.56 $\pm$ 0.16
RealMLP 9.57 $\pm$ 0.29 9.44 $\pm$ 0.23 9.22 $\pm$ 0.21
TabM 9.51 $\pm$ 0.21 9.31 $\pm$ 0.21 9.24 $\pm$ 0.20
MNCA 9.56 $\pm$ 0.16 9.50 $\pm$ 0.14 9.28 $\pm$ 0.19
TabPFNv2 - - -
TabDPT 9.08 $\pm$ 0.20 - -
TabICL - - -
Linear 17.43 $\pm$ 0.10 17.42 $\pm$ 0.09 17.27 $\pm$ 0.10
KNN 12.38 $\pm$ 0.14 10.85 $\pm$ 0.27 10.63 $\pm$ 0.21
AutoGluon - - 9.22 $\pm$ 0.16
taiwanese_bankruptcy_prediction (AUC $\uparrow$)
Default Tuned Tuned + Ens.
RF 0.933 $\pm$ 0.011 0.932 $\pm$ 0.011 0.932 $\pm$ 0.012
ExtraTrees 0.937 $\pm$ 0.011 0.937 $\pm$ 0.008 0.938 $\pm$ 0.008
XGBoost 0.941 $\pm$ 0.008 0.943 $\pm$ 0.008 0.944 $\pm$ 0.008
LightGBM 0.938 $\pm$ 0.008 0.943 $\pm$ 0.008 0.944 $\pm$ 0.007
CatBoost 0.944 $\pm$ 0.006 0.944 $\pm$ 0.007 0.943 $\pm$ 0.006
EBM 0.942 $\pm$ 0.005 0.940 $\pm$ 0.005 0.941 $\pm$ 0.004
FastAIMLP 0.914 $\pm$ 0.022 0.923 $\pm$ 0.009 0.927 $\pm$ 0.010
TorchMLP 0.925 $\pm$ 0.009 0.940 $\pm$ 0.007 0.943 $\pm$ 0.007
RealMLP 0.941 $\pm$ 0.006 0.941 $\pm$ 0.007 0.945 $\pm$ 0.006
TabM 0.941 $\pm$ 0.004 0.942 $\pm$ 0.004 0.943 $\pm$ 0.004
MNCA 0.940 $\pm$ 0.006 0.939 $\pm$ 0.010 0.932 $\pm$ 0.013
TabPFNv2 0.942 $\pm$ 0.005 0.943 $\pm$ 0.007 0.945 $\pm$ 0.008
TabDPT 0.937 $\pm$ 0.008 - -
TabICL 0.944 $\pm$ 0.006 - -
Linear 0.936 $\pm$ 0.005 0.936 $\pm$ 0.005 0.936 $\pm$ 0.005
KNN 0.594 $\pm$ 0.018 0.678 $\pm$ 0.029 0.681 $\pm$ 0.027
AutoGluon - - 0.946 $\pm$ 0.006
website_phishing (logloss $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.312 $\pm$ 0.037 0.262 $\pm$ 0.019 0.257 $\pm$ 0.019
ExtraTrees 0.312 $\pm$ 0.036 0.258 $\pm$ 0.020 0.250 $\pm$ 0.019
XGBoost 0.260 $\pm$ 0.027 0.251 $\pm$ 0.022 0.251 $\pm$ 0.023
LightGBM 0.255 $\pm$ 0.021 0.249 $\pm$ 0.021 0.247 $\pm$ 0.021
CatBoost 0.252 $\pm$ 0.022 0.239 $\pm$ 0.022 0.239 $\pm$ 0.021
EBM 0.357 $\pm$ 0.021 0.357 $\pm$ 0.020 0.357 $\pm$ 0.020
FastAIMLP 0.334 $\pm$ 0.023 0.245 $\pm$ 0.022 0.241 $\pm$ 0.021
TorchMLP 0.289 $\pm$ 0.035 0.237 $\pm$ 0.020 0.230 $\pm$ 0.019
RealMLP 0.256 $\pm$ 0.026 0.236 $\pm$ 0.026 0.232 $\pm$ 0.021
TabM 0.246 $\pm$ 0.026 0.238 $\pm$ 0.027 0.236 $\pm$ 0.025
MNCA 0.261 $\pm$ 0.023 0.243 $\pm$ 0.025 0.239 $\pm$ 0.022
TabPFNv2 0.229 $\pm$ 0.025 0.223 $\pm$ 0.027 0.222 $\pm$ 0.025
TabDPT 0.228 $\pm$ 0.027 - -
TabICL 0.228 $\pm$ 0.026 - -
Linear 0.360 $\pm$ 0.021 0.361 $\pm$ 0.022 0.358 $\pm$ 0.023
KNN 0.312 $\pm$ 0.037 0.312 $\pm$ 0.037 0.312 $\pm$ 0.037
AutoGluon - - 0.233 $\pm$ 0.021
wine_quality (rmse $\downarrow$)
Default Tuned Tuned + Ens.
RF 0.620 $\pm$ 0.021 0.616 $\pm$ 0.021 0.611 $\pm$ 0.021
ExtraTrees 0.613 $\pm$ 0.021 0.606 $\pm$ 0.025 0.603 $\pm$ 0.021
XGBoost 0.621 $\pm$ 0.021 0.610 $\pm$ 0.019 0.610 $\pm$ 0.020
LightGBM 0.628 $\pm$ 0.019 0.607 $\pm$ 0.019 0.608 $\pm$ 0.019
CatBoost 0.621 $\pm$ 0.019 0.605 $\pm$ 0.020 0.605 $\pm$ 0.020
EBM 0.679 $\pm$ 0.015 0.678 $\pm$ 0.016 0.675 $\pm$ 0.016
FastAIMLP 0.669 $\pm$ 0.019 0.668 $\pm$ 0.018 0.657 $\pm$ 0.019
TorchMLP 0.691 $\pm$ 0.019 0.653 $\pm$ 0.018 0.650 $\pm$ 0.015
RealMLP 0.625 $\pm$ 0.018 0.618 $\pm$ 0.019 0.603 $\pm$ 0.020
TabM 0.637 $\pm$ 0.020 0.620 $\pm$ 0.020 0.612 $\pm$ 0.020
MNCA 0.610 $\pm$ 0.020 0.606 $\pm$ 0.019 0.601 $\pm$ 0.020
TabPFNv2 0.692 $\pm$ 0.012 0.639 $\pm$ 0.017 0.610 $\pm$ 0.020
TabDPT 0.590 $\pm$ 0.018 - -
TabICL - - -
Linear 0.731 $\pm$ 0.017 0.731 $\pm$ 0.017 0.730 $\pm$ 0.016
KNN 0.809 $\pm$ 0.018 0.689 $\pm$ 0.020 0.687 $\pm$ 0.020
AutoGluon - - 0.599 $\pm$ 0.020